### v1

# Double Robust Variance Estimation

Abstract
Doubly robust estimators have gained popularity in the field of causal inference due to their ability to provide consistent point estimates when either an outcome or exposure model is correctly specified. However, the influence function based variance estimator frequently used with doubly robust estimators is only consistent when both the outcome and exposure models are correctly specified. Here, use of M-estimation and the empirical sandwich variance estimator for doubly robust point and variance estimation is demonstrated. Simulation studies illustrate the properties of the influence function based and empirical sandwich variance estimators. Estimators are applied to data from the Improving Pregnancy Outcomes with Progesterone (IPOP) trial to estimate the effect of maternal anemia on birth weight among women with HIV. In the example, birth weights if all women had anemia were estimated to be lower than birth weights if no women had anemia, though estimates were imprecise. Variance estimates were more stable under varying model specifications for the empirical sandwich variance estimator than the influence function based variance estimator.
Keywords causal inference; double robustness; empirical sandwich variance; augmented inverse probability weighting; M-estimation.
1 Introduction
Two common approaches for estimating the causal effect of a point exposure or treatment with observational studies are methods based on inverse probability of treatment weighting (IPTW) and the g-formula. IPTW estimators (Robins,, 1998; Robins et al.,, 2000; Hernán et al.,, 2000) rely on estimation of the probability of exposure given a set of covariates, i.e., the propensity score. Alternatively, the g-formula (Robins,, 1986) is a direct standardization approach where the outcome is modeled as a function of the exposure and covariates. When estimated nonparametrically, IPTW and g-formula estimators are equivalent (Hernán and Robins,, 2020). However, in high dimensional settings (e.g., settings with continuous covariates), analysts typically resort to parametric models for the exposure and outcome, and the IPTW and g-formula estimators are no longer equivalent due to differing parametric constraints. Both approaches provide consistent estimators of the average causal effect (ACE) if the corresponding model is correctly specified. Because the functional forms of these “nuisance” models are typically unknown, correct specification remains an obstacle to valid effect estimation.
Augmented inverse probability weighted (AIPW) estimators and targeted maximum likelihood estimators (TMLE) combine estimates from the exposure and outcome models in such a way that resulting estimators are “doubly robust.” That is, they are consistent when either the exposure or outcome model is correctly specified, but not necessarily both (Robins et al.,, 2007; Bang and Robins,, 2005; Funk et al.,, 2011; Kang and Schafer,, 2007; Hernán and Robins,, 2020; van der Laan and Rubin,, 2006). In other words, doubly robust methods give the analyst two opportunities for correct model specification. Two common AIPW estimators are an estimator that incorporates estimated propensity scores from the exposure model and pseudo-outcomes from the outcome model (i.e., the classic AIPW) (Hernán and Robins,, 2020; Lunceford and Davidian,, 2004) and a g-formula type estimator where the outcome model is weighted by the IPTW (i.e., the weighted regression AIPW) (Kang and Schafer,, 2007; Robins et al.,, 2007; Vansteelandt and Keiding,, 2011; Gabriel et al.,, 2023). TMLE is a variation of the weighted regression AIPW estimator that instead combines IPTW and the outcome model through a “targeting” model fit using the outcome predictions and weights (van der Laan and Rubin,, 2006). Both the weighted regression AIPW estimator and TMLE respect the parameter space of the estimand, while the classic AIPW can produce estimates outside of the range of the parameter space (Gabriel et al.,, 2023; Gruber and Van Der Laan,, 2012).
In addition to the double robustness property, doubly robust estimators also offer precision advantages over IPTW estimators. When both nuisance models are correctly specified, the three doubly robust estimators previously discussed are semiparametric efficient, having the smallest asymptotic variance of all IPTW estimators (Robins et al.,, 1994; van der Laan and Rubin,, 2006; Gabriel et al.,, 2023). This means that even in settings where the analyst is confident in the specification of one of the models (e.g., the exposure mechanism is well known), doubly robust estimators can offer improved efficiency relative to IPTW estimators, provided that the outcome model is not grossly misspecified. Given these potential precision advantages and that the true specification of both models is typically unknown, doubly robust approaches have gained popularity (Sullivan et al.,, 2022; Josey et al.,, 2023; Lee et al.,, 2019; Zhong et al.,, 2018; Gerdts et al.,, 2022; Chua et al.,, 2022; Hoffman et al.,, 2022; Yazdi et al.,, 2021).
Despite adoption of doubly robust estimators, there is a major limitation with the corresponding influence function based variance estimator commonly used in applications and provided by software. Namely, this variance estimator is only consistent if both the propensity and outcome models are correctly specified (Daniel,, 2014). This limitation has been highlighted previously (Funk et al.,, 2011; Daniel,, 2014), yet the influence function based variance estimator is still commonly used. A systematic review published in October 2023 found that 67% of papers published using TMLE did not indicate how variances were estimated. Of the 33% that did report the method, the majority (73%) relied on the influence function based variance estimator (Smith et al.,, 2023). Widespread application of the influence function based variance estimator is aided by available software, as it is used for confidence interval (CI) construction in the R AIPW (Zhong et al.,, 2021) and tmle (Gruber and Van Der Laan,, 2012) packages as well as the zEpid (Zivich et al., 022a, ) Python package.
The influence function based variance estimator’s reliance on correctly specifying both exposure and outcome models is particularly concerning in instances where researchers understand either the exposure or outcome mechanism well, but not both. If the outcome process is well known but the exposure process is not well understood, leading to correct specification of the outcome model but misspecification of the propensity model, the influence function based variance estimator is inconsistent (Laan and Robins,, 2003; Muñoz and Van Der Laan,, 2012). In this setting, consistent point estimators may be paired with CIs that do not achieve the nominal level of coverage. Alternatively, if the exposure process is well understood but the outcome process is not, leading to correct specification of the propensity model but misspecification of the outcome model, the influence function based variance estimator is conservative (Muñoz and Van Der Laan,, 2012). Precision advantages of doubly robust estimators may be lost with application of a conservative variance estimator. For example, we aim to estimate the effect of maternal anemia on birth weight among women with HIV. While most causes of anemia among women with HIV are well understood, doubly robust estimators may be preferred over IPTW estimators due to efficiency advantages. Modeling the relationship between anemia and birth weight among women with HIV is challenging due to complex dependencies between anemia and other drivers of birth weight. Furthermore, low birth weight may result from two separate pathways - preterm birth or fetal growth restrictions - each of which is a complex syndrome with several causes. For these reasons, the outcome model may be susceptible to misspecification, thus application of the influence function based variance estimator may lead to conservative variance estimates in this study.
Alternative, but less commonly used, variance estimation approaches are themselves doubly robust, providing consistent estimates of the variance when only one of the two nuisance models is correctly specified. Two such approaches are the nonparametric bootstrap and the empirical sandwich variance estimator based on M-estimation theory. The nonparametric bootstrap, while doubly robust (Funk et al.,, 2011), is computationally burdensome. For large datasets, this computational burden may preclude its use. M-estimation (Stefanski and Boos,, 2002) has been proposed for use with doubly robust estimators (Lunceford and Davidian,, 2004; Daniel,, 2014), but we are not aware of prior work highlighting the doubly robust property of the empirical sandwich variance estimator. Furthermore, the recent systematic review (Smith et al.,, 2023) did not mention application of the empirical sandwich variance estimator in any of the papers evaluated. This suggests that the doubly robust property of the empirical sandwich variance estimator is not well known.
This paper demonstrates how M-estimation can be applied for AIPW estimators and TMLE to obtain doubly robust point and variance estimators. In Section 2, three common doubly robust estimators are described along with two approaches for inference: the influence function based variance estimator and M-estimation paired with the empirical sandwich variance estimator. Estimators are applied to data from the Improving Pregnancy Outcomes with Progesterone (IPOP) trial to estimate the average causal effect of anemia on birth weight in Section 3. The doubly robust property of the empirical sandwich variance estimator is demonstrated in a simulation study and is contrasted with the performance of the influence function based variance estimator in Section 4. The benefits and limitations of each variance estimation approach are discussed in Section 5. R code for implementing each of the doubly robust estimators is provided on GitHub.
2 Methods
2.1 Preliminaries
[…]
2.2 Estimators
Three doubly robust approaches for estimating the $ACE$ are considered: the classic AIPW estimator, the weighted regression AIPW estimator, and TMLE. All three methods combine estimates from a model for the probability of exposure conditional on covariates (propensity model) and a model for the outcome conditional on exposure and covariates (outcome model). The approach for modeling the propensity score is common across the three methods, while there are slight differences in fitting the outcome model. For all three approaches, the propensity score for each participant, $e_{i}=\Pr(X_{i}=1\mid Z_{i})$, is estimated. In practice, often the model $\mbox{logit}(e_{i})=G_{i}^{T}\alpha$ is fit, where $G_{i}=g(Z_{i})$ is a vector of predictors for participant $i$ for some user-specified function $g$ of $Z_{i}$ and $\alpha$ is the vector of regression coefficients from the propensity model. The predicted propensity score for each participant is calculated as $\hat{e}_{i}=e(Z_{i},\hat{\alpha})=\mbox{logit}^{-1}(G_{i}^{T}\hat{\alpha})$ where $\hat{\alpha}$ is the maximum likelihood estimate (MLE) of $\alpha$. The IPTW is then estimated for each participant as $\hat{W}_{i}=X_{i}\hat{e}_{i}^{-1}+(1-X_{i})(1-\hat{e}_{i})^{-1}$.
2.2.1 Classic AIPW estimator
For the classic AIPW estimator, the $Y\mid Z,X$ model is fit and used to predict pseudo-outcomes for each participant under both exposures. That is, $a_{x}(Z_{i},\hat{\gamma})=\hat{E}(Y_{i}\mid Z_{i},X_{i}=x)$ for $x\in\{0,1\}$, where $\hat{\gamma}$ are the MLEs of the parameters from the assumed outcome model. Then, the classic AIPW estimator for the $ACE$ is
$\widehat{DR}_{C}=\hat{\mu}_{C}^{1}-\hat{\mu}_{C}^{0}$ (1)
where $\hat{\mu}_{C}^{1}=n^{-1}\sum_{i=1}^{n}\hat{e}_{i}^{-1}\{X_{i}Y_{i}-(X_{i}-\hat{e}_{i})a_{1}(Z_{i},\hat{\gamma})\}$ and $\hat{\mu}_{C}^{0}=n^{-1}\sum_{i=1}^{n}(1-\hat{e}_{i})^{-1}\{(1-X_{i})Y_{i}+(X_{i}-\hat{e}_{i})a_{0}(Z_{i},\hat{\gamma})\}$. The estimator (1) was originally proposed by Robins et al., (1994) and was further examined in simulation studies, such as Lunceford and Davidian, (2004), Kang and Schafer, (2007), and Funk et al., (2011).
2.2.2 Weighted regression AIPW estimator
An alternative AIPW estimator incorporates the estimated propensity scores when fitting the outcome regression model. The parameters of the $Y\mid Z,X$ model are estimated with IPTW-weighted maximum likelihood estimation. Pseudo-outcomes are obtained for each participant under both exposures, i.e., $b_{x}(Z_{i},\hat{\beta})=\hat{E}(Y_{i}\mid Z_{i},X_{i}=x)$ for $x\in\{0,1\}$, where $\hat{\beta}$ are the MLEs from the weighted outcome model. The weighted regression AIPW estimator is
$\widehat{DR}_{WR}=\hat{\mu}_{WR}^{1}-\hat{\mu}_{WR}^{0}$ (2)
where $\hat{\mu}_{WR}^{x}=n^{-1}\sum_{i=1}^{n}b_{x}(Z_{i},\hat{\beta})$ for $x\in\{0,1\}$. The weighted regression AIPW estimator (2) has been evaluated previously (Kang and Schafer,, 2007; Robins et al.,, 2007; Vansteelandt and Keiding,, 2011; Gabriel et al.,, 2023) and is expected to be more stable than (1) when IPTWs are extreme due to bounding within the parameter space (Kang and Schafer,, 2007; Robins et al.,, 2007; Vansteelandt et al.,, 2012).
2.2.3 TMLE
TMLE is an alternative doubly robust estimator that consists of a two-stage process. In the first stage, $Y\mid Z,X$ is modeled as in Section 2.2.1, and $a_{x}(Z_{i},\hat{\gamma})$ are calculated for $x\in\{0,1\}$. In the second stage, referred to as the targeting stage (van der Laan and Rubin,, 2006), corrections to outcome regression predictions from stage one are made that incorporate the estimated propensity scores. Different methods have been proposed for fitting the targeting model; here, the weighted regression approach is considered, as it is thought to handle random positivity violations better (van der Laan et al.,, 2011). Specifically, the models $\mbox{logit}(Y_{i})=\eta_{x}+\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}$ are fit for $x\in\{0,1\}$. Parameters $\eta_{x}$ are estimated using weighted maximum-likelihood, with weights $(1-X)\hat{W}_{i}$ and $X\hat{W}_{i}$ for the $x=0$ and $x=1$ targeting model, respectively. In the targeting models, logit-transformed pseudo-outcomes ($\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}$) for each participant are included as offset terms, and MLEs $\hat{\eta}_{0}$ and $\hat{\eta}_{1}$ can be thought of as corrections to these pseudo-outcomes (Zivich et al., 022b, ). The stage two pseudo-outcomes are computed as $c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})=\mbox{expit}[\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}+\hat{\eta}_{x}]$. The TMLE estimator of $ACE$ is
$\widehat{DR}_{TMLE}=\hat{\mu}_{TMLE}^{1}-\hat{\mu}_{TMLE}^{0}$ (3)
where $\hat{\mu}_{TMLE}^{x}=n^{-1}\sum_{i=1}^{n}c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})$ for $x\in\{0,1\}$. Note that for continuous outcomes, such as birth weight considered in Section 3, the outcome is scaled prior to implementing TMLE, i.e., $Y_{i}^{*}=(Y_{i}-a)/(b-a)$ is defined for each participant, where $(a,b)$ are the bounds of $Y$. Then, $Y_{i}$ is replaced with $Y_{i}^{*}$ in the steps above, and $c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})=\mbox{expit}[\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}+\hat{\eta}_{x}](b-a)+a$ re-scales the outcome back to the original bounds of $Y$.
TMLE methods were originally proposed by van der Laan and Rubin, (2006) and have been extended and applied in a number of settings (Stitelman et al.,, 2012; Balzer et al.,, 2023; van der Laan et al.,, 2018; van der Laan and Gruber,, 2012). When parametric models are used in both stages of estimation, (3) is expected to be similar to (2) (Tran et al.,, 2019).
2.3 Variance Estimation
Variance estimation for doubly robust estimators (1), (2), and (3) is not straightforward, because the variance of each estimator depends on estimating the parameters of the nuisance models. Here, we consider two methods for variance estimation: the commonly used influence function based variance estimator and the empirical sandwich variance estimator. Both variance estimators can be used to construct Wald-typed CIs for the $ACE$.
2.3.1 Influence function based variance estimator
Estimators (1), (2), and (3) are suggested by the efficient influence function (Luque-Fernandez et al.,, 2018; Gabriel et al.,, 2023). The commonly used influence function based variance estimator has the form (Glynn and Quinn,, 2010; Gruber and Van Der Laan,, 2012)
$\hat{V}(\widehat{DR})=n^{-2}\sum_{i=1}^{n}\hat{I}_{i}^{2}$ (4)
where $\hat{I}_{i}=\hat{e}_{i}^{-1}X_{i}Y_{i}-(1-\hat{e}_{i})^{-1}(1-X_{i})Y_{i}-\{\hat{e}_{i}^{-1}(1-\hat{e}_{i})^{-1}(X_{i}-\hat{e}_{i})\}\{(1-\hat{e}_{i})\hat{Y}_{i}^{1}+\hat{e}_{i}\hat{Y}_{i}^{0}\}-\widehat{DR}$. Here, pseudo-outcomes $\hat{Y}_{i}^{x}$ equal $a_{x}(Z_{i},\hat{\gamma})$, $b_{x}(Z_{i},\hat{\beta})$, and $c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})$ for the classic AIPW, weighted regression AIPW, and TMLE methods, respectively, and $\widehat{DR}$ equals (1), (2), and (3) for the three corresponding estimators.
The estimators (4) are consistent for the variance of (1), (2), and (3) when both propensity score and outcome models are correctly specified (Daniel,, 2014; Gruber and Van Der Laan,, 2012). When one of the two nuisance models is misspecified, the influence function based variance estimator is not consistent. As noted in the introduction, the estimator (4) is conservative when the weight model is correctly specified but the outcome model is misspecified (Muñoz and Van Der Laan,, 2012), and may be conservative or anti-conservative when the outcome model is correctly specified but the weight model is misspecified (Laan and Robins,, 2003; Muñoz and Van Der Laan,, 2012). In empirical studies, the variance estimator (4) for the classic AIPW estimator (1) was found to be unstable for small sample sizes ($n<1000$), even under correct model specification (Funk et al.,, 2011). Similar instability has been reported for (4) with TMLE ($n=100$) (van der Laan and Gruber,, 2011).
2.3.2 Empirical sandwich variance estimator
An alternative to the influence function based variance estimator (4) is the empirical sandwich variance estimator. The estimators (1), (2), and (3) can each be expressed as the solution to a set of unbiased estimating equations. The M-estimator for each set of estimating equations, $\hat{\theta}$, is the solution (for $\theta$) to $\sum_{i=1}^{n}\psi_{q}(O_{i},\theta)=0$, where $\psi_{q}(O_{i},\theta)$ are the estimating functions corresponding to each of the three estimators $q\in\{1,2,3\}$ with parameter vector $\theta$. Parameter vectors and estimating functions for (1), (2), and (3) can be specified as follows.
Classic AIPW estimator
For the classic AIPW estimator, $\theta=[\alpha,\gamma,\mu_{C}^{1},\mu_{C}^{0},DR_{C}]$. The estimating function $\psi_{1}$ is specified as
$\psi_{1}(O_{i};\theta)=\begin{bmatrix}\psi_{\alpha}\\ \psi_{\gamma}\\ \psi_{\mu_{C1}}\\ \psi_{\mu_{C0}}\\ \psi_{C}\end{bmatrix}=\begin{bmatrix}\{X_{i}-\mbox{expit}(G_{i}^{T}\alpha)\}G_{i}\\ \{Y_{i}-\phi^{-1}(H_{i}^{T}\gamma)\}H_{i}\\ \{X_{i}Y_{i}-(X_{i}-e_{i})a_{1}(Z_{i},\gamma)\}e_{i}^{-1}-\mu_{C}^{1}\\ \{(1-X_{i})Y_{i}+(X_{i}-e_{i})a_{0}(Z_{i},\gamma)\}(1-e_{i})^{-1}-\mu_{C}^{0}\\ \mu_{C}^{1}-\mu_{C}^{0}-DR_{C}\end{bmatrix}$
where $\psi_{\alpha}$ is the vector of score functions from the propensity score model, $\psi_{\gamma}$ is the vector of score functions for the outcome model (with link function $\phi$), and $H_{i}=h(X_{i},Z_{i})$ is a vector of predictors for the outcome model for participant $i$ for some user-specified function $h$ of $X_{i}$ and $Z_{i}$. The estimating functions $\psi_{\mu_{C1}}$ and $\psi_{\mu_{C0}}$ are for the causal means $\mu_{C}^{1}$ and $\mu_{C}^{0}$, respectively, and $\psi_{C}$ is a delta method transformation of the causal means for estimation of the $ACE$ using the classic AIPW, i.e., $\widehat{DR}_{C}$. Shook-Sa et al., (2023) demonstrate that this set of estimating equations is unbiased when at least one of the models is correctly specified.
Weighted regression AIPW estimator
For the weighted regression AIPW estimator, $\theta=[\alpha,\beta,\mu_{WR}^{1},\mu_{WR}^{0},DR_{WR}]$. The estimating function $\psi_{2}$ is specified as
$\psi_{2}(O_{i};\theta)=\begin{bmatrix}\psi_{\alpha}\\ \psi_{\beta}\\ \psi_{\mu_{WR1}}\\ \psi_{\mu_{WR0}}\\ \psi_{WR}\end{bmatrix}=\begin{bmatrix}\{X_{i}-\mbox{expit}(G_{i}^{T}\alpha)\}G_{i}\\ W_{i}\{Y_{i}-\phi^{-1}(H_{i}^{T}\beta)\}H_{i}\\ b_{1}(Z_{i},\beta)-\mu_{WR}^{1}\\ b_{0}(Z_{i},\beta)-\mu_{WR}^{0}\\ \mu_{WR}^{1}-\mu_{WR}^{0}-DR_{WR}\end{bmatrix}$
where $\psi_{\beta}$ is the vector of IPTW-weighted score functions for the outcome model. The estimating functions $\psi_{\mu_{WR1}}$ and $\psi_{\mu_{WR0}}$ are for the causal means $\mu_{WR}^{1}$ and $\mu_{WR}^{0}$, respectively, and $\psi_{WR}$ is a delta method transformation of the causal means for estimation of the $ACE$ using the weighted AIPW estimator $\widehat{DR}_{WR}$. Gabriel et al., (2023) demonstrate unbiasedness of this set of estimating equations when at least one of the nuisance models is correctly specified.
TMLE
With TMLE, $\theta=[\alpha,\gamma,\eta_{1},\eta_{0},\mu_{TMLE}^{1},\mu_{TMLE}^{0},DR_{TMLE}]$. For continuous outcomes, the estimating function $\psi_{3}$ is specified as
$\psi_{3}(O_{i};\theta)=\begin{bmatrix}\psi_{\alpha}\\ \psi_{\gamma}\\ \psi_{t1}\\ \psi_{t0}\\ \psi_{\mu_{TMLE1}}\\ \psi_{\mu_{TMLE0}}\\ \psi_{TMLE}\end{bmatrix}=\begin{bmatrix}\{X_{i}-\mbox{expit}(G_{i}^{T}\alpha)\}G_{i}\\ \{Y^{*}_{i}-\phi^{-1}(H_{i}^{T}\gamma)\}H_{i}\\ X_{i}e_{i}^{-1}(Y^{*}_{i}-\mbox{expit}[\eta_{1}+\mbox{logit}\{a_{1}(Z_{i},\gamma)\}])\\ (1-X_{i})(1-e_{i})^{-1}(Y^{*}_{i}-\mbox{expit}[\eta_{0}+\mbox{logit}\{a_{0}(Z_{i},\gamma)\}])\\ c_{1}(O_{i},\gamma,\eta_{1})-\mu_{TMLE}^{1}\\ c_{0}(O_{i},\gamma,\eta_{0})-\mu_{TMLE}^{0}\\ \mu_{TMLE}^{1}-\mu_{TMLE}^{0}-DR_{TMLE}\end{bmatrix}$
where $\psi_{\gamma}$ is the vector of score functions for the scaled outcome model. The estimating functions $\psi_{t1}$ and $\psi_{t0}$ are for the two targeting models, while $\psi_{\mu_{TMLE1}}$ and $\psi_{\mu_{TMLE0}}$ are for the causal means $\mu_{TMLE}^{1}$ and $\mu_{TMLE}^{0}$, respectively. Note $c_{x}(O_{i},\gamma,\eta_{x})=\mbox{expit}[\mbox{logit}\{a_{x}(Z_{i},\gamma)\}+\eta_{x}](b-a)+a$ are the pseudo-outcomes following the targeting step, as outlined in Section 2.2.3. Finally, $\psi_{TMLE}$ is a delta method transformation of the causal means for estimation of the $ACE$ using $\widehat{DR}_{TMLE}$. Unbiasedness of $\psi_{3}(O_{i};\theta)$ when at least one nuisance model is correctly specified is demonstrated in the Appendix.
[…]
3 Example: Improving Pregnancy Outcomes with Progesterone
Prior research has shown links between maternal anemia and low birth weight (Azizah et al.,, 2022; Figueiredo et al.,, 2018). The relationship between anemia and birth weight is particularly important to understand for women living with HIV, as they experience anemia more frequently than women without HIV (Levine et al.,, 2001) and are more prone to adverse birth outcomes (Tukei et al.,, 2021). The effect of maternal anemia on birth weight among women living with HIV was estimated with data from the IPOP trial.
The IPOP trial was a randomized, double-blind, placebo-controlled trial conducted in Lusaka, Zambia. Participants were pregnant women aged 18 or older living with HIV on (or intending to initiate) antiretroviral therapy (ART) who had viable singleton pregnancies at less than 24 weeks of gestation at baseline. The 800 enrolled participants were randomized with equal probability to either a placebo or 17 alpha-hydroxyprogesterone caproate (17P) treatment to evaluate the efficacy of 17P for prevention of preterm birth. Additional details about the IPOP trial have been published previously (Price et al.,, 2021). While IPOP was designed to examine the effect of 17P on preterm birth or stillbirth, covariates collected during the trial allow for examination of other causal questions, assuming that the identification assumptions hold.
The three doubly robust estimators described in Section 2.2 were applied to IPOP data for estimation of the $ACE$ of maternal anemia on birth weight. The exposure, maternal anemia, was defined as having a baseline hemoglobin level below 10.5 g/dL. The IPOP data were limited to the 782 participants (98%) with no prior preterm births and with measured birth weights and baseline hemoglobin values. Note 14 participants experienced stillbirth. Birth weights for these participants were included in the analysis; we return to this issue in the discussion. It was assumed that potential outcomes are independent of the exposure conditional on the following baseline covariates: trimester (1 or 2), maternal age group (18-24, 25-29, 30-34, 35-39, 40+), number of prior births (0, 1, 2, 3, 4+), alcohol use during pregnancy (yes or no), ART use (yes or no), and maternal height (cm) (Table 1). At baseline, no IPOP participants reported tobacco or drug use during pregnancy, so these covariates were not included. Primary analyses were conducted using R and were independently replicated in Python.
As discussed in the introduction, application of doubly robust estimators for estimating the $ACE$ of maternal anemia on birth weight may offer precision advantages over IPTW estimators, but the correct model specifications are unknown. To demonstrate differences in the influence function based variance estimator and the empirical sandwich variance estimator presented in Section 2.3, each of the doubly robust estimators were applied with three sets of model specifications. First, the covariates in Table 1 were included in both propensity and outcome models. Then, a naive outcome model was fit that included only the exposure variable (anemia status), while the propensity model included the full set of covariates. Under this specification, the outcome model was reasonably thought to be misspecified due to its simplicity. Finally, a naive propensity model was fit that included only an intercept term, while the outcome model included the full set of covariates. Due to the simplicity of the propensity model, it was reasonably thought to be misspecified. For all estimators, maternal height was modeled using restricted cubic splines with four knots placed at the 5th, 35th, 65th, and 95th percentiles (Perperoglou et al.,, 2019). The empirical sandwich variance estimator was computed using geex in R (Saul and Hudgens,, 2020) and delicatessen in Python (Zivich et al., 022c, ). Confidence intervals for the $ACE$ were constructed based on the influence function based variance estimator and the empirical sandwich variance estimator. Estimates for each of the three methods and model specifications along with corresponding 95% CIs are presented in Table 2, with 95% CI half-widths compared in Figure 1.
IPOP participants with and without anemia at baseline had similar self-reported alcohol use during pregnancy and similar age and height distributions. Participants without anemia were more likely to have one or more prior births, be in their first trimester, and be on ART. The three doubly robust methods provided similar estimates of the $ACE$ for all model specifications examined, with average birth weights if all women had anemia were approximately 40g lower than birth weights if no women had anemia, though all 95% CIs included zero.
Precision estimates based on the empirical sandwich variance estimator and the influence function based variance estimator were similar when the full covariate set was included in both the outcome and propensity models. Under the naive outcome model, standard errors for the influence function based variance estimator were larger than those for the empirical sandwich variance estimator, resulting in wider CIs. When the propensity model was naive, standard errors for the influence function based variance estimator were smaller than those for the empirical sandwich variance estimator, resulting in narrower CIs. Note there was more fluctuation in 95% CI half-widths across model specifications for the influence function based variance estimator than the empirical sandwich variance estimator (Figure 1).
Table 1: Baseline characteristics of IPOP participants by anemia status
No anemia $n=664$ Anemia $n=118$
Age category 18-24 156 (23%) 28 (24%)
25-29 190 (29%) 33 (28%)
30-34 192 (29%) 31 (26%)
35-39 100 (15%) 21 (18%)
40+ 26 (4%) 5 (4%)
Number of prior births 0 125 (19%) 30 (25%)
1 160 (24%) 25 (21%)
2 157 (24%) 28 (24%)
3 125 (19%) 18 (15%)
4+ 97 (15%) 17 (14%)
First trimester 120 (18%) 12 (10%)
ART use 648 (98%) 109 (92%)
Alcohol use during pregnancy 64 (10%) 9 (8%)
Height Median (Q1,Q3) 155 (150, 160) 154 (150, 159)
Mean (SD) 155 (8) 153 (8)
Min, Max 131, 185 132, 170
  • Note: Anemia defined as baseline hemoglobin below 10.5 g/dL
Table 2: Estimated effect of maternal anemia on birth weight by model specification and estimator
ACE ES-SE ES-95% CI IF-SE IF-95% CI
Full Covariate Set
Classic AIPW -37 56 (-147, 73) 58 (-151, 76)
Weighted regression AIPW -41 56 (-151, 69) 56 (-151, 68)
TMLE -37 56 (-147, 73) 58 (-151, 76)
Naive Outcome Model
Classic AIPW -36 57 (-148, 76) 61 (-156, 84)
Weighted regression AIPW -36 57 (-148, 77) 61 (-156, 84)
TMLE -36 57 (-148, 77) 61 (-156, 84)
Naive Propensity Model
Classic AIPW -41 58 (-154, 73) 56 (-150, 69)
Weighted regression AIPW -47 57 (-159, 65) 54 (-153, 59)
TMLE -40 58 (-154, 73) 56 (-150, 69)
  • Note: SE=standard error, ES=empirical sandwich, IF=influence function, CI=confidence interval. The naive outcome model included only the exposure, and the naive propensity model included only an intercept.
Figure 1: 95% confidence interval (CI) half-widths by model specification and estimator. FC=Full Covariate Set, NO=Naive Outcome Model, NP=Naive Propensity Model
4 Simulation Study
A simulation study was conducted to compare the empirical properties of the influence function based variance estimator and the empirical sandwich variance estimator in conjunction with each of the three doubly robust methods discussed in Section 2.
4.1 Simulation Setup
Simulations were designed based on the example in Section 3, which aimed to estimate the effect of maternal anemia on birth weight among women with HIV. The primary simulations were conducted with $n=800$ participants, similar to the sample size in the example. Results for a sample size of $n=2000$ are included in the Supporting Information. The simulation study was conducted in R with code provided on GitHub.
Three covariates ($Z_{1}$, $Z_{2}$, $Z_{3}$), the exposure ($X$), and potential outcomes ($Y^{0}$ and $Y^{1}$) were simulated. The covariate $Z_{1}$ was distributed Normal with mean $155$ and standard deviation $7.6$. Two binary covariates $Z_{2}$ and $Z_{3}$ were simulated from Bernoulli distributions with means $0.25$ and $0.75$, respectively. The exposure $X$ was simulated from a Bernoulli distribution with mean $\mbox{expit}(15-0.1Z_{1}+2.5Z_{2}-1Z_{3}-0.02Z_{1}Z_{2}+0.005Z_{1}Z_{3})$. Potential outcomes $Y^{0}$ and $Y^{1}$ under exposure and no exposure, respectively, were simulated from a normal distribution with mean $E(Y^{x})=1000+11.5Z_{1}+100Z_{2}-15Z_{1}Z_{2}+25x-5.5xZ_{1}-30xZ_{2}+20xZ_{1}Z_{2}$ and standard deviation $\sigma=400$ for $x\in\{0,1\}$. Under this data generating mechanism, the $ACE$ was approximately $-60$. Additional simulations were conducted with $\sigma=200$ with results in the Supporting Information.
The estimators $\widehat{DR}_{C}$, $\widehat{DR}_{WR}$, and $\widehat{DR}_{TMLE}$ were applied to 5000 simulated samples to estimate the $ACE$. Estimators were computed four ways: (1) with both propensity and outcome models correctly specified, (2) with propensity models correctly specified but outcome models misspecied, (3) with outcome models correctly specified but propensity models misspecified, and (4) with both models misspecified. Misspecified propensity models included only an intercept and a linear term for $(Z_{1}-155)^{2}$, and misspecified outcome models included only an intercept and linear terms for $X$ and $(Z_{1}-155)^{2}$.
[…]
4.2 Simulation Results
Both point and variance estimators performed as expected in simulations. The results of the simulation study for $n=800$ and $\sigma=400$ are presented in Table 3 and Figure 2. When at least one model was correctly specified, the classic AIPW, weighted regression AIPW, and TMLE displayed minimal bias, but all were substantially biased when both nuisance models were misspecified. Note bias was further reduced under correct specification of at least one model for $n=2000$ (Table 4). Under correct specification of both nuisance models, the empirical sandwich and influence function based variance estimators tracked closely with the ESE, resulting in SERs close to one and CIs attaining the nominal level of coverage. When both models were misspecified, both variance estimators tracked with the ESE, but bias was substantial, resulting in CIs with below nominal coverage.
Differences between the variance estimators are apparent when examining the simulation study results for scenarios where one nuisance model was correctly specified and the other was misspecified. When either the outcome model or the propensity model was misspecified, CIs based on the empirical sandwich variance estimator attained the nominal level of coverage, and estimated standard errors tracked closely with the ESE. In Figure 2, note variance ratios for the empirical sandwich estimator are closely clustered around one for these scenarios. In contrast, the influence function based variance estimator was empirically biased when either nuisance model was misspecified. As expected, it tended to overestimate the variance when the outcome model was misspecified, leading to estimated standard errors that tracked above the ESE, SERs above one, and conservative CIs that exceeded the nominal level of coverage. When the propensity model was misspecified but the outcome model was correctly specified, the influence function based variance estimator underestimated the variance, resulting in estimated standard errors below the ESE and SERs below one. The biases observed for the influence function based variance estimator were also present when $n=2000$ (Table 4 and Figure 3).
In the $\sigma=200$ scenarios, the empirical sandwich variance estimator performance was similar to the $\sigma=400$ scenario, but the performance of the influence function based variance estimator varied (Tables 5-6 and Figures 4-5). Specifically, when the propensity model was misspecified but the outcome model was correctly specified, the influence function based variance estimator demonstrated minimal empirical bias, resulting in SERs close to one.
In summary, the simulation study demonstrated the theoretical properties explained in Section 2. That is, the empirical sandwich variance estimator was consistent when at least one of the two nuisance models was correctly specified, and performed well in simulations. The influence function based variance estimator is not consistent under misspecification of either nuisance model. In line with expectations, the influence function based variance estimator was conservative when the outcome model was misspecified. For this data generating mechanism, it was anti-conservative when the propensity model was misspecified for $\sigma=400$ and exhibited minimal bias when $\sigma=200$.
Figure 2: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, n=800, $\sigma=400$, 5000 simulations. Black squares denote the mean variance ratio (=SER). Results exclude one simulation where models failed to converge. Note the $0.3\%$ of correct model specification simulations and $4.5\%$ of misspecified outcome model simulations where the ratio was above 1.2 are not displayed.
Table 3: Simulation summary results, $n=800$, $\sigma=400$, $5000$ simulations. Bias, Relative Bias, ASE, ESE, SER, and 95% CI coverage calculated for the ACE.
Scenario Estimator Bias Relative Bias (%) ESE ASE, ES SER, ES Cov, ES (%) ASE, IF SER, IF Cov, IF (%)
CS Classic 0.4 0.6 58.4 58.0 0.99 95 58.1 1.00 95
WR 0.4 0.6 58.3 57.7 0.99 95 57.6 0.99 95
TMLE 0.4 0.6 58.3 57.9 0.99 95 58.1 1.00 95
MO Classic -0.4 0.6 60.0 59.2 0.99 95 64.5 1.07 97
WR -1.7 2.8 59.2 58.4 0.99 95 62.4 1.05 96
TMLE -0.5 0.8 59.7 58.9 0.99 95 63.0 1.06 96
MW Classic 0.3 0.5 57.8 57.6 1.00 95 55.8 0.97 94
WR 0.2 0.4 57.7 57.6 1.00 95 55.8 0.97 94
TMLE 0.3 0.5 57.8 57.6 1.00 95 55.8 0.97 94
MB Classic -23.8 39.8 57.0 56.8 1.00 92 56.8 1.00 92
WR -23.8 39.8 57.0 56.8 1.00 92 56.8 1.00 92
TMLE -23.8 39.8 57.0 56.8 1.00 92 56.8 1.00 92
  • ASE=average estimated standard error; ESE=empirical standard error; SER=standard error ratio (ASE/ESE); Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MW=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%. Results exclude one simulation where models failed to converge.
5 Discussion
Doubly robust estimators have increased in popularity in the field of causal inference due to their ability to provide consistent point estimates when either an outcome or propensity model is correctly specified (Smith et al.,, 2023). In this paper, the commonly-used influence function based variance estimator is compared with the empirical sandwich variance estimator in conjunction with three doubly robust estimators: the classic AIPW estimator, the weighted regression AIPW estimator, and TMLE. These three estimators can be specified as solutions to sets of unbiased estimating equations, and thus their variances can be consistently estimated by the empirical sandwich variance estimator. Because the estimating equations remain unbiased when either the outcome or propensity model is correctly specified, the empirical sandwich variance estimator is a doubly robust variance estimator. In contrast, the influence function based variance estimator is consistent only when both outcome and propensity models are correctly specified. This discrepancy may be of particular importance in settings where the functional form of one of the nuisance models is known but the other is not. For example, with randomized trials, randomization to treatment means that the exposure mechanism is known. Doubly robust estimators can offer improved efficiency in this setting, but some of that efficiency gain may not be realized with application of the influence function based variance estimator, which is conservative if the outcome model is misspecified.
While consideration in this paper was limited to finite-dimensional parametric modeling approaches, machine learning approaches can also be used with doubly robust estimators (Kennedy et al.,, 2016; van der Laan and Gruber,, 2011, 2012; van der Laan et al.,, 2018; Balzer et al.,, 2023). Machine learning approaches allow for more complex functional forms for continuous covariates than fully parametric approaches and include higher order interaction terms that may not be included in investigator-specified parametric models, often leading to estimators that are more robust to model misspecification (Zivich et al.,, 2022). However, convergence of machine learning algorithms is typically slower and consistent variance estimation is more challenging than with parametric modeling approaches. Machine learning methods are typically not compatible with the estimating equations approach discussed in this paper, though alternative methods have been developed for doubly robust variance estimation in this context for TMLE estimators (Benkeser et al.,, 2017).
In the example, the effect of anemia on birth weight among women with HIV was estimated by applying the three doubly robust methods to data from the IPOP randomized trial. Average birth weights if all women had anemia were estimated to be approximately 40g lower than birth weights if no women had anemia, though all 95% CIs included zero. The analysis included 14 participants who experienced stillbirth. Because stillbirth is a competing outcome for live birth weight, future work could reexamine this problem using alternative approaches for competing events.
Use of the empirical sandwich variance estimator in conjunction with doubly robust estimators is not new, though its doubly robust property has not been emphasized. Lunceford and Davidian, (2004) discuss use of the empirical sandwich variance estimator for estimating the variance of the classic AIPW estimator and note that it tends to be more stable than the influence function based variance estimator. However, based on the systematic review discussed in the introduction (Smith et al.,, 2023), the influence function based variance estimator is frequently applied and the empirical sandwich variance estimator does not appear to be widely used with TMLE. While some existing software packages compute empirical sandwich variance estimators for doubly robust estimators (e.g., the “causaltrt” procedure in SAS and the “dr” procedure in Stata), other popular software packages compute the influence function based variance estimator but not the empirical sandwich variance estimator (e.g., the AIPW and tmle packages in R and the zEpid package in Python). We provide example R and Python code for estimating the variance of each of the three doubly robust estimators with the empirical sandwich variance estimator, which we hope will allow for easier implementation of doubly robust point and variance estimation.
Supplementary Material
R and Python code for computing the different estimators along with the corresponding standard error estimators is available at https://github.com/bonnieshook/DoublyRobustVariance.
Acknowledgements
This work was supported by US National Institutes of Health under award numbers R01 AI157758, R01 AI085073, P30 AI050410, and K01 AI177102. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health. The authors thank Bradley Saul at the University of North Carolina at Chapel Hill for coding support.
Appendix
Consistency proof of TMLE estimator under correct specification of either the outcome or propensity model
Recall the TMLE estimating function $\psi_{3}$ is specified as
$\psi_{3}(O_{i};\theta)=\begin{bmatrix}\psi_{\alpha}\\ \psi_{\gamma}\\ \psi_{t1}\\ \psi_{t0}\\ \psi_{\mu_{TMLE1}}\\ \psi_{\mu_{TMLE0}}\\ \psi_{TMLE}\end{bmatrix}=\begin{bmatrix}\{X_{i}-\mbox{expit}(G_{i}^{T}\alpha)\}G_{i}\\ \{Y^{*}_{i}-\phi^{-1}(H_{i}^{T}\gamma)\}H_{i}\\ X_{i}e_{i}^{-1}(Y^{*}_{i}-\mbox{expit}[\eta_{1}+\mbox{logit}\{a_{1}(Z_{i},\gamma)\}])\\ (1-X_{i})(1-e_{i})^{-1}(Y^{*}_{i}-\mbox{expit}[\eta_{0}+\mbox{logit}\{a_{0}(Z_{i},\gamma)\}])\\ c_{1}(O_{i},\gamma,\eta_{1})-\mu_{TMLE}^{1}\\ c_{0}(O_{i},\gamma,\eta_{0})-\mu_{TMLE}^{0}\\ \mu_{TMLE}^{1}-\mu_{TMLE}^{0}-DR_{TMLE}\end{bmatrix}$
where $\theta=[\alpha,\gamma,\eta_{1},\eta_{0},\mu_{TMLE}^{1},\mu_{TMLE}^{0},DR_{TMLE}]$ is the vector of parameters.
Let
$\tilde{\theta}=[\tilde{\alpha},\tilde{\gamma},\tilde{\eta}_{1},\tilde{\eta}_{0},\tilde{\mu}_{TMLE}^{1},\tilde{\mu}_{TMLE}^{0},\widetilde{DR}_{TMLE}]$ be the solution of $E\{\psi_{3}(O_{i};\theta)\}=0$ and $\widehat{DR}_{TMLE}$ be the proposed TMLE estimator of ACE, obtained by solving the estimating equation $n^{-1}\sum_{i}\psi_{3}(O_{i};\theta)=0$. Then, if either the outcome model or the propensity score model is correctly specified, $\sqrt{n}(\widehat{DR}_{TMLE}-ACE)\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,\sigma^{2}(\tilde{\theta}))$, where $\sigma^{2}(\tilde{\theta})$ is the bottom right element of $V({\tilde{\theta}})=V({\theta})|_{\theta=\tilde{\theta}}$, and $V({\theta})=A(\theta)^{-1}B(\theta)A(\theta)^{-T}$, with $A(\theta)=E\{-\partial\psi_{3}(O_{i};\theta)/\partial\theta\}$, $B(\theta)=E\{\psi_{3}(O_{i};\theta)\psi_{3}(O_{i};\theta)^{\top}\}$, which is a doubly robust variance estimator of $\widehat{DR}_{TMLE}$.
Proof.
Let $\hat{\theta}$ be the solution of the estimating equation $n^{-1}\sum_{i}\psi_{3}(O_{i};\theta)=0$. Under suitable regularity conditions (Stefanski and Boos,, 2002), $\sqrt{n}(\hat{\theta}-\tilde{\theta})\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,V(\tilde{\theta}))$. Note the final element of $\hat{\theta}$ is $\widehat{DR}_{TMLE}$ and therefore $\sqrt{n}(\widehat{DR}_{TMLE}-\widetilde{DR}_{TMLE})\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,\sigma^{2}(\tilde{\theta}))$. It remains to show that if either the outcome model or the propensity score model is correctly specified, $\widetilde{DR}_{TMLE}=ACE$.
First, let $\theta_{0}=[\alpha_{0},\gamma_{0},\eta_{10},\eta_{00},\mu_{TMLE0}^{1},\mu_{TMLE0}^{0},DR_{TMLE0}]$ be the true parameter values from the data generating process. Then, $P(X_{i}=1|Z_{i})=e(Z_{i},\alpha_{0})=\mbox{expit}(G_{i}^{T}\alpha_{0})$ and $E(Y_{i}^{*}|Z_{i},X_{i})=\phi^{-1}(H_{i}^{T}\gamma_{0})=X_{i}a_{1}(Z_{i},\gamma_{0})+(1-X_{i})a_{0}(Z_{i},\gamma_{0})$.
Now assume the outcome model is correctly specified, i.e., $\tilde{\gamma}=\gamma_{0}$ and $E(Y_{i}^{*}|Z_{i},X_{i})=X_{i}a_{1}(Z_{i},\gamma_{0})+(1-X_{i})a_{0}(Z_{i},\gamma_{0})$. From $E\{\psi_{t1}(O_{i};\tilde{\theta})\}=0$,
$0=$ $E\{\psi_{t1}(O_{i};\tilde{\theta})\}$
$=$ $E[E\{\psi_{t1}(O_{i};\tilde{\theta})|X_{i}\}]$
$=$ $E[E\{X_{i}e(Z_{i},\tilde{\alpha})^{-1}(Y_{i}^{*}-\mbox{expit}[\tilde{\eta}_{1}+\mbox{logit}\{a_{1}(Z_{i},\gamma_{0})\}])|X_{i}\}]$
$=$ $E\{e(Z_{i},\tilde{\alpha})^{-1}(Y_{i}^{*}-\mbox{expit}[\tilde{\eta}_{1}+\mbox{logit}\{a_{1}(Z_{i},\gamma_{0})\}])|X_{i}=1\}P(X_{i}=1)$
$=$ $E[E\{e(Z_{i},\tilde{\alpha})^{-1}(Y_{i}^{*}-\mbox{expit}[\tilde{\eta}_{1}+\mbox{logit}\{a_{1}(Z_{i},\gamma_{0})\}])|Z_{i},X_{i}=1\}|X_{i}=1]P(X_{i}=1)$
$=$ $E[e(Z_{i},\tilde{\alpha})^{-1}\{E(Y_{i}^{*}|Z_{i},X_{i}=1)-\mbox{expit}[\tilde{\eta}_{1}+\mbox{logit}\{a_{1}(Z_{i},\gamma_{0})\}]\}|X_{i}=1]P(X_{i}=1)$
$=$ $E[e(Z_{i},\tilde{\alpha})^{-1}\{a_{1}(Z_{i},\gamma_{0})-\mbox{expit}[\tilde{\eta}_{1}+\mbox{logit}\{a_{1}(Z_{i},\gamma_{0})\}]\}|X_{i}=1]P(X_{i}=1)$
$=$ $E[e(Z_{i},\tilde{\alpha})^{-1}F(\tilde{\eta}_{1})|X_{i}=1]P(X_{i}=1)$
where $F(t)=\mbox{expit}(d)-\mbox{expit}(t+d)$ and $d=\mbox{logit}\{a_{1}(Z_{i},\gamma_{0})\}$. Since the function $F(t)$ is strictly decreasing in $t$, it can be shown that the last line in the above is strictly decreasing in $\tilde{\eta}_{1}$ and attains 0 only at $\tilde{\eta}_{1}=0$. Therefore, $\tilde{\eta}_{1}=0$. Then, from $E\{\psi_{\mu_{TMLE1}}(O_{i};\tilde{\theta})\}=0$,
$\tilde{\mu}_{TMLE}^{1}=$ $E(Y_{i,TMLE}^{1})$
$=$ $E\{\mbox{expit}[\tilde{\eta}_{1}+\mbox{logit}\{a_{1}(Z_{i},\gamma_{0})\}](b-a)+a\}$
$=$ $E\{a_{1}(Z_{i},\gamma_{0})\}(b-a)+a$
$=$ $E[E\{Y_{i}^{*}|Z_{i},X_{i}=1\}](b-a)+a$
$=$ $E[E\{(Y_{i}-a)/(b-a)|Z_{i},X_{i}=1\}](b-a)+a$
$=$ $E\{E(Y_{i}|Z_{i},X_{i}=1)\}$
$=$ $\mu^{1}.$
Similarly, $\tilde{\mu}_{TMLE}^{0}=\mu^{0}$.
Next, assume the propensity score model is correctly specified, i.e., $\tilde{\alpha}=\alpha_{0}$ and $P(X_{i}=1|Z_{i})=e(Z_{i},\alpha_{0})$. From $E\{\psi_{t1}(O_{i};\tilde{\theta})\}=0$,
$0=$ $E\{\psi_{t1}(O_{i};\tilde{\theta})\}$
$=$ $E[E\{\psi_{t1}(O_{i};\tilde{\theta})|Z_{i}\}]$
$=$ $E[E\{X_{i}e(Z_{i},\alpha_{0})^{-1}\{(Y_{i}^{1}-a)/(b-a)-\mbox{expit}[\tilde{\eta}_{1}+\mbox{logit}\{a_{1}(Z_{i},\tilde{\gamma})\}])\}|Z_{i}\}]$
$=$ $E(E(X_{i}|Z_{i})e(Z_{i},\alpha_{0})^{-1}[E\{(Y_{i}^{1}-a)/(b-a)|Z_{i}\}-\mbox{expit}[\tilde{\eta}_{1}+\mbox{logit}\{a_{1}(Z_{i},\tilde{\gamma})\}]])$
$=$ $E[E\{(Y_{i}^{1}-a)/(b-a)|Z_{i}\}]-E(\mbox{expit}[\tilde{\eta}_{1}+\mbox{logit}\{a_{1}(Z_{i},\tilde{\gamma})\}])$
$=$ $(\mu^{1}-a)/(b-a)-E(\mbox{expit}[\tilde{\eta}_{1}+\mbox{logit}\{a_{1}(Z_{i},\tilde{\gamma})\}]).$
Therefore, $E(\mbox{expit}[\tilde{\eta}_{1}+\mbox{logit}\{a_{1}(Z_{i},\tilde{\gamma})\}])=(\mu^{1}-a)/(b-a)$. Then, from $E\{\psi_{\mu_{TMLE1}}(O_{i};\tilde{\theta})\}=0$,
$\tilde{\mu}_{TMLE}^{1}=$ $E\{c_{x}(O_{i},\gamma,\eta_{x})\}$
$=$ $E\{\mbox{expit}[\tilde{\eta}_{1}+\mbox{logit}\{a_{1}(Z_{i},\tilde{\gamma})\}](b-a)+a\}$
$=$ $\mu^{1}.$
Similarly, $\tilde{\mu}_{TMLE}^{0}=\mu^{0}$.
In conclusion, if either the outcome model or the propensity score model is correctly specified, $\widetilde{DR}_{TMLE}=\tilde{\mu}_{TMLE}^{1}-\tilde{\mu}_{TMLE}^{0}=\mu^{1}-\mu^{0}=ACE$ which proves the stated result.
Table 4: Simulation summary results, $n=2000$, $\sigma=400$, $5000$ simulations. Bias, Relative Bias, ASE, ESE, SER, and 95% CI coverage calculated for the ACE.
Scenario Estimator Bias Relative Bias (%) ESE ASE, ES SER, ES Cov, ES (%) ASE, IF SER, IF Cov, IF (%)
CS Classic -0.1 0.2 36.6 36.7 1.00 95 36.8 1.00 95
WR -0.1 0.2 36.6 36.6 1.00 95 36.6 1.00 95
TMLE -0.1 0.2 36.6 36.7 1.00 95 36.8 1.00 95
MO Classic -0.3 0.6 37.2 37.4 1.01 95 40.6 1.09 97
WR -0.9 1.6 37.0 37.1 1.00 95 39.6 1.07 96
TMLE -0.4 0.6 37.1 37.3 1.01 95 39.8 1.07 96
MW Classic -0.2 0.3 36.2 36.4 1.01 95 35.3 0.98 94
WR -0.2 0.3 36.2 36.4 1.01 95 35.3 0.98 95
TMLE -0.2 0.3 36.2 36.4 1.01 95 35.3 0.98 94
MB Classic -23.8 39.8 35.9 35.9 1.00 90 35.9 1.00 90
WR -23.8 39.8 35.9 35.9 1.00 90 35.9 1.00 90
TMLE -23.8 39.8 35.9 35.9 1.00 90 35.9 1.00 90
  • ASE=average estimated standard error; ESE=empirical standard error; SER=standard error ratio (ASE/ESE); Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MW=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
Table 5: Simulation summary results, $n=800$, $\sigma=200$, $5000$ simulations. Bias, Relative Bias, ASE, ESE, SER, and 95% CI coverage calculated for the ACE.
Scenario Estimator Bias Relative Bias (%) ESE ASE, ES SER, ES Cov, ES (%) ASE, IF SER, IF Cov, IF (%)
CS Classic 0.3 0.6 50.2 50.0 1.00 95 50.0 1.00 95
WR 0.3 0.5 50.1 49.9 1.00 95 49.9 1.00 95
TMLE 0.3 0.5 50.2 50.0 1.00 95 50.0 1.00 95
MO Classic -0.4 0.6 52.0 51.5 0.99 95 57.2 1.10 97
WR -1.6 2.7 51.3 50.7 0.99 95 55.1 1.07 96
TMLE -0.5 0.8 51.6 51.1 0.99 95 55.6 1.08 96
MW Classic 0.3 0.5 50.0 49.9 1.00 95 49.4 0.99 95
WR 0.3 0.4 50.0 49.9 1.00 95 49.4 0.99 95
TMLE 0.3 0.5 50.0 49.9 1.00 95 49.4 0.99 95
MB Classic -23.6 39.4 50.7 50.4 0.99 92 50.4 0.99 92
WR -23.6 39.4 50.7 50.4 0.99 92 50.4 0.99 92
TMLE -23.5 39.3 50.7 50.4 0.99 92 50.4 0.99 92
  • ASE=average estimated standard error; ESE=empirical standard error; SER=standard error ratio (ASE/ESE); Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MW=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%. Results exclude 14 simulations where models did not converge.
Table 6: Simulation summary results, $n=2000$, $\sigma=200$, $5000$ simulations. Bias, Relative Bias, ASE, ESE, SER, and 95% CI coverage calculated for the ACE.
Scenario Estimator Bias Relative Bias (%) ESE ASE, ES SER, ES Cov, ES (%) ASE, IF SER, IF Cov, IF (%)
CS Classic 0.0 0.0 31.6 31.6 1.00 95 31.7 1.00 95
WR 0.0 0.1 31.5 31.6 1.00 95 31.6 1.00 95
TMLE 0.0 0.0 31.6 31.6 1.00 95 31.7 1.00 95
MO Classic -0.2 0.4 32.3 32.4 1.00 95 36.0 1.11 97
WR -0.8 1.4 32.1 32.1 1.00 95 34.9 1.09 97
TMLE -0.3 0.4 32.1 32.3 1.00 95 35.1 1.09 97
MW Classic 0.0 0.0 31.4 31.6 1.00 95 31.2 0.99 95
WR 0.0 0.0 31.4 31.6 1.00 95 31.2 0.99 95
TMLE 0.0 0.0 31.4 31.6 1.00 95 31.2 0.99 95
MB Classic -23.6 39.5 32.0 31.9 1.00 89 31.9 1.00 89
WR -23.6 39.5 32.0 31.9 1.00 89 31.9 1.00 89
TMLE -23.6 39.5 32.0 31.9 1.00 89 31.9 1.00 89
  • ASE=average estimated standard error; ESE=empirical standard error; SER=standard error ratio (ASE/ESE); Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MW=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
Figure 3: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, n=2000, $\sigma=400$, 5000 simulations. Black squares denote the mean variance ratio (=SER). Note the $0.08\%$ of correct model specification simulations and $3.4\%$ of misspecified outcome model simulations where the ratio was above 1.2 are not displayed.
Figure 4: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, n=800, $\sigma=200$, 5000 simulations. Black squares denote the mean variance ratio (=SER). Results exclude 14 simulations where models did not converge. Note the $0.06\%$ of correct model specification simulations, $5.2\%$ of misspecified outcome model simulations, and $0.01\%$ of simulations where both models were misspecified where the ratio was above 1.2 or below 0.8 are not displayed.
Figure 5: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, n=2000, $\sigma=200$, 5000 simulations. Black squares denote the mean variance ratio (=SER). Note the $3.9\%$ of misspecified outcome model simulations where the ratio was above 1.2 are not displayed.

### v2

@@
-Double Robust Variance Estimation
+Double Robust Variance Estimation with Parametric Working Models
 Abstract
 Doubly robust estimators have gained popularity in the field of causal inference due to their ability to provide consistent point estimates when either an outcome or exposure model is correctly specified.
-However, the influence function based variance estimator frequently used with doubly robust estimators is only consistent when both the outcome and exposure models are correctly specified.
-Here, use of M-estimation and the empirical sandwich variance estimator for doubly robust point and variance estimation is demonstrated.
-Simulation studies illustrate the properties of the influence function based and empirical sandwich variance estimators.
-Estimators are applied to data from the Improving Pregnancy Outcomes with Progesterone (IPOP) trial to estimate the effect of maternal anemia on birth weight among women with HIV.
-In the example, birth weights if all women had anemia were estimated to be lower than birth weights if no women had anemia, though estimates were imprecise.
-Variance estimates were more stable under varying model specifications for the empirical sandwich variance estimator than the influence function based variance estimator.
-Keywords causal inference; double robustness; empirical sandwich variance; augmented inverse probability weighting; M-estimation.
+However, for nonrandomized exposures the influence function based variance estimator frequently used with doubly robust estimators of the average causal effect is only consistent when both working models (i.e., outcome and exposure models) are correctly specified.
+Here, the empirical sandwich variance estimator and the nonparametric bootstrap are demonstrated to be doubly robust variance estimators.
+That is, they are expected to provide valid estimates of the variance leading to nominal confidence interval coverage when only one working model is correctly specified.
+Simulation studies illustrate the properties of the influence function based, empirical sandwich, and nonparametric bootstrap variance estimators in the setting where parametric working models are assumed.
+Estimators are applied to data from the Improving Pregnancy Outcomes with Progesterone (IPOP) study to estimate the effect of maternal anemia on birth weight among women with HIV.
+Keywords augmented inverse probability weighting; causal inference; double robustness; empirical sandwich variance; M-estimation.
 1 Introduction
-Two common approaches for estimating the causal effect of a point exposure or treatment with observational studies are methods based on inverse probability of treatment weighting (IPTW) and the g-formula.
-IPTW estimators (Robins,, 1998; Robins et al.,, 2000; Hernán et al.,, 2000) rely on estimation of the probability of exposure given a set of covariates, i.e., the propensity score.
-Alternatively, the g-formula (Robins,, 1986) is a direct standardization approach where the outcome is modeled as a function of the exposure and covariates.
-When estimated nonparametrically, IPTW and g-formula estimators are equivalent (Hernán and Robins,, 2020).
-However, in high dimensional settings (e.g., settings with continuous covariates), analysts typically resort to parametric models for the exposure and outcome, and the IPTW and g-formula estimators are no longer equivalent due to differing parametric constraints.
-Both approaches provide consistent estimators of the average causal effect (ACE) if the corresponding model is correctly specified.
-Because the functional forms of these “nuisance” models are typically unknown, correct specification remains an obstacle to valid effect estimation.
-Augmented inverse probability weighted (AIPW) estimators and targeted maximum likelihood estimators (TMLE) combine estimates from the exposure and outcome models in such a way that resulting estimators are “doubly robust.” That is, they are consistent when either the exposure or outcome model is correctly specified, but not necessarily both (Robins et al.,, 2007; Bang and Robins,, 2005; Funk et al.,, 2011; Kang and Schafer,, 2007; Hernán and Robins,, 2020; van der Laan and Rubin,, 2006).
-In other words, doubly robust methods give the analyst two opportunities for correct model specification.
-Two common AIPW estimators are an estimator that incorporates estimated propensity scores from the exposure model and pseudo-outcomes from the outcome model (i.e., the classic AIPW) (Hernán and Robins,, 2020; Lunceford and Davidian,, 2004) and a g-formula type estimator where the outcome model is weighted by the IPTW (i.e., the weighted regression AIPW) (Kang and Schafer,, 2007; Robins et al.,, 2007; Vansteelandt and Keiding,, 2011; Gabriel et al.,, 2023).
-TMLE is a variation of the weighted regression AIPW estimator that instead combines IPTW and the outcome model through a “targeting” model fit using the outcome predictions and weights (van der Laan and Rubin,, 2006).
-Both the weighted regression AIPW estimator and TMLE respect the parameter space of the estimand, while the classic AIPW can produce estimates outside of the range of the parameter space (Gabriel et al.,, 2023; Gruber and Van Der Laan,, 2012).
-In addition to the double robustness property, doubly robust estimators also offer precision advantages over IPTW estimators.
-When both nuisance models are correctly specified, the three doubly robust estimators previously discussed are semiparametric efficient, having the smallest asymptotic variance of all IPTW estimators (Robins et al.,, 1994; van der Laan and Rubin,, 2006; Gabriel et al.,, 2023).
-This means that even in settings where the analyst is confident in the specification of one of the models (e.g., the exposure mechanism is well known), doubly robust estimators can offer improved efficiency relative to IPTW estimators, provided that the outcome model is not grossly misspecified.
-Given these potential precision advantages and that the true specification of both models is typically unknown, doubly robust approaches have gained popularity (Sullivan et al.,, 2022; Josey et al.,, 2023; Lee et al.,, 2019; Zhong et al.,, 2018; Gerdts et al.,, 2022; Chua et al.,, 2022; Hoffman et al.,, 2022; Yazdi et al.,, 2021).
-Despite adoption of doubly robust estimators, there is a major limitation with the corresponding influence function based variance estimator commonly used in applications and provided by software.
-Namely, this variance estimator is only consistent if both the propensity and outcome models are correctly specified (Daniel,, 2014).
+Two common approaches for estimating causal effects of point exposures with observational study data are inverse probability of treatment weighting (IPTW) and g-computation.
+IPTW estimators rely on estimation of the probability of exposure given a set of covariates, i.e., the propensity score.
+Alternatively, g-computation is a direct standardization approach where the outcome is modeled as a function of the exposure and covariates.
+Both approaches provide consistent estimators of the average causal effect (ACE) if the corresponding model is correctly specified, but these estimators are generally not equivalent due to differing parametric constraints imposed by modeling.
+As the functional forms of these working, or nuisance, models are typically unknown, correct specification remains an obstacle to valid effect estimation.
+Augmented inverse probability weighted (AIPW) estimators and targeted maximum likelihood estimators (TMLE) combine estimates from the working models in such a way that resulting estimators are “doubly robust.” That is, they are consistent when either the exposure or outcome model is correctly specified, but not necessarily both (Robins et al.,, 2007; Funk et al.,, 2011; Kang and Schafer,, 2007; van der Laan and Rubin,, 2006).
+Two common variations of AIPW estimators are an estimator that incorporates estimated propensity scores from the exposure model and pseudo-outcomes from the outcome model (i.e., the classic AIPW) and a g-computation type estimator where the outcome model is weighted by the IPTW (i.e., the weighted regression AIPW).
+TMLE is a variation of the weighted regression AIPW estimator that instead combines IPTW and the outcome model through a “targeting” model fit using the outcome predictions and weights.
+While primarily developed with machine learning approaches for estimation of nuisance parameters (van der Laan and Rubin,, 2006), TMLE can also be applied with parametric working models.
+Doubly robust estimators also offer precision advantages over IPTW estimators.
+When both working models are correctly specified, the three doubly robust estimators previously discussed are semiparametric efficient, having the smallest asymptotic variance of all IPTW estimators (Robins et al.,, 1994; van der Laan and Rubin,, 2006; Gabriel et al.,, 2023).
+Given these potential precision advantages and that the true specification of both models is typically unknown, doubly robust approaches have gained popularity.
+The influence function based variance estimator commonly used with doubly robust methods has a major limitation.
+When estimating the average causal effect from observational data, this variance estimator is only consistent if both working models are correctly specified, i.e., this variance estimator is not doubly robust.
 This limitation has been highlighted previously (Funk et al.,, 2011; Daniel,, 2014), yet the influence function based variance estimator is still commonly used.
 A systematic review published in October 2023 found that 67% of papers published using TMLE did not indicate how variances were estimated.
-Of the 33% that did report the method, the majority (73%) relied on the influence function based variance estimator (Smith et al.,, 2023).
-Widespread application of the influence function based variance estimator is aided by available software, as it is used for confidence interval (CI) construction in the R AIPW (Zhong et al.,, 2021) and tmle (Gruber and Van Der Laan,, 2012) packages as well as the zEpid (Zivich et al., 022a, ) Python package.
-The influence function based variance estimator’s reliance on correctly specifying both exposure and outcome models is particularly concerning in instances where researchers understand either the exposure or outcome mechanism well, but not both.
-If the outcome process is well known but the exposure process is not well understood, leading to correct specification of the outcome model but misspecification of the propensity model, the influence function based variance estimator is inconsistent (Laan and Robins,, 2003; Muñoz and Van Der Laan,, 2012).
+Of the 33% that did report the method, the majority (72%) relied on the influence function based variance estimator (Smith et al.,, 2023).
+Widespread application of the influence function based variance estimator is aided by available software, as it is used for confidence interval (CI) construction in the R AIPW and tmle packages and the zEpid Python package.
+When estimating the average causal effect, performance of the influence function based variance estimator depends on which working model is misspecified.
+When the outcome model is correctly specified but the propensity model is misspecified, the influence function based variance estimator is inconsistent (Muñoz and Van Der Laan,, 2012).
 In this setting, consistent point estimators may be paired with CIs that do not achieve the nominal level of coverage.
-Alternatively, if the exposure process is well understood but the outcome process is not, leading to correct specification of the propensity model but misspecification of the outcome model, the influence function based variance estimator is conservative (Muñoz and Van Der Laan,, 2012).
+If the propensity model is correctly specified but the outcome model is misspecified, the influence function based variance estimator is conservative (Muñoz and Van Der Laan,, 2012).
 Precision advantages of doubly robust estimators may be lost with application of a conservative variance estimator.
-For example, we aim to estimate the effect of maternal anemia on birth weight among women with HIV.
-While most causes of anemia among women with HIV are well understood, doubly robust estimators may be preferred over IPTW estimators due to efficiency advantages.
-Modeling the relationship between anemia and birth weight among women with HIV is challenging due to complex dependencies between anemia and other drivers of birth weight.
-Furthermore, low birth weight may result from two separate pathways - preterm birth or fetal growth restrictions - each of which is a complex syndrome with several causes.
-For these reasons, the outcome model may be susceptible to misspecification, thus application of the influence function based variance estimator may lead to conservative variance estimates in this study.
-Alternative, but less commonly used, variance estimation approaches are themselves doubly robust, providing consistent estimates of the variance when only one of the two nuisance models is correctly specified.
+These properties of the influence function based variance estimator do not necessarily apply for other estimands and in the randomized setting (see Section 5).
+Alternative, but less commonly used, variance estimation approaches are doubly robust, as they are expected to provide valid estimates of the variance leading to nominal CI coverage when only one working model is correctly specified.
 Two such approaches are the nonparametric bootstrap and the empirical sandwich variance estimator based on M-estimation theory.
-The nonparametric bootstrap, while doubly robust (Funk et al.,, 2011), is computationally burdensome.
-For large datasets, this computational burden may preclude its use.
-M-estimation (Stefanski and Boos,, 2002) has been proposed for use with doubly robust estimators (Lunceford and Davidian,, 2004; Daniel,, 2014), but we are not aware of prior work highlighting the doubly robust property of the empirical sandwich variance estimator.
-Furthermore, the recent systematic review (Smith et al.,, 2023) did not mention application of the empirical sandwich variance estimator in any of the papers evaluated.
-This suggests that the doubly robust property of the empirical sandwich variance estimator is not well known.
-This paper demonstrates how M-estimation can be applied for AIPW estimators and TMLE to obtain doubly robust point and variance estimators.
-In Section 2, three common doubly robust estimators are described along with two approaches for inference: the influence function based variance estimator and M-estimation paired with the empirical sandwich variance estimator.
-Estimators are applied to data from the Improving Pregnancy Outcomes with Progesterone (IPOP) trial to estimate the average causal effect of anemia on birth weight in Section 3.
-The doubly robust property of the empirical sandwich variance estimator is demonstrated in a simulation study and is contrasted with the performance of the influence function based variance estimator in Section 4.
+The doubly robust property of the nonparametric bootstrap has been illustrated previously (Funk et al.,, 2011).
+M-estimation has been proposed for use with doubly robust estimators (Lunceford and Davidian,, 2004; Daniel,, 2014), but we are not aware of prior work highlighting the doubly robust property of the empirical sandwich variance estimator.
+The M-estimation approach is equivalent to using the correct influence function with correction terms for estimation of nuisance parameters (Stefanski and Boos,, 2002).
+The recent systematic review (Smith et al.,, 2023) did not mention application of the empirical sandwich variance estimator in any reviewed papers.
+M-estimation is appropriate with finite-dimension parametric working models and is generally not compatible with many machine learning approaches.
+SuperLearner was used in the majority of studies considered in the systematic review (62%), and the empirical sandwich variance estimator would be inappropriate for these studies.
+However, logistic regression was used in 20% of studies considered, and an additional 18% of studies did not indicate what methods were used for nuisance parameter estimation.
+The lack of use of the empirical sandwich variance estimator in these settings suggests that the doubly robust property of the empirical sandwich variance estimator is not well known.
+Among published studies that reported the variance estimation approach, 21% applied the bootstrap.
+However, knowledge of its doubly robust property may or may not be widespread.
+We aim to estimate the effect of maternal anemia on birth weight among women with HIV.
+Prior research has shown links between maternal anemia and low birth weight (Azizah et al.,, 2022), and women with HIV experience anemia more frequently than women without HIV (Levine et al.,, 2001).
+This effect is estimated with data from approximately 800 participants from the Improving Pregnancy Outcomes with Progesterone (IPOP) study.
+IPOP was a randomized controlled trial conducted in Lusaka, Zambia to evaluate the effectiveness of a therapeutic for prevention of preterm birth and stillbirth among women with HIV.
+Additional details about the IPOP trial have been published previously (Price et al.,, 2021).
+Covariates collected during the trial allow for examination of the relationship between anemia and birth weight.
+Doubly robust estimators may be preferred over IPTW or g-computation approaches due to uncertainty in working model specifications and potential efficiency advantages.
+This paper demonstrates how M-estimation and the nonparametric bootstrap can be applied with AIPW estimators and TMLE to obtain doubly robust point and variance estimators with parametric working models.
+In Section 2, three common doubly robust estimators are described along with three approaches for inference: the influence function based variance estimator, the empirical sandwich variance estimator, and the nonparametric bootstrap.
+Estimators are applied to IPOP data in Section 3.
+In Section 4, the doubly robust properties of the empirical sandwich variance estimator and the nonparametric bootstrap are demonstrated in a simulation study and are contrasted with the performance of the influence function based variance estimator.
 The benefits and limitations of each variance estimation approach are discussed in Section 5.
-R code for implementing each of the doubly robust estimators is provided on GitHub.
 2 Methods
@@
 2.2 Estimators
-Three doubly robust approaches for estimating the $ACE$ are considered: the classic AIPW estimator, the weighted regression AIPW estimator, and TMLE.
-All three methods combine estimates from a model for the probability of exposure conditional on covariates (propensity model) and a model for the outcome conditional on exposure and covariates (outcome model).
-The approach for modeling the propensity score is common across the three methods, while there are slight differences in fitting the outcome model.
-For all three approaches, the propensity score for each participant, $e_{i}=\Pr(X_{i}=1\mid Z_{i})$, is estimated.
-In practice, often the model $\mbox{logit}(e_{i})=G_{i}^{T}\alpha$ is fit, where $G_{i}=g(Z_{i})$ is a vector of predictors for participant $i$ for some user-specified function $g$ of $Z_{i}$ and $\alpha$ is the vector of regression coefficients from the propensity model.
-The predicted propensity score for each participant is calculated as $\hat{e}_{i}=e(Z_{i},\hat{\alpha})=\mbox{logit}^{-1}(G_{i}^{T}\hat{\alpha})$ where $\hat{\alpha}$ is the maximum likelihood estimate (MLE) of $\alpha$.
+Three doubly robust estimators of $ACE$ are considered, each of which combines estimates from the propensity and outcome models.
+The approach for modeling the propensity score is common across the methods, while there are slight differences in fitting the outcome model.
+In practice, the propensity score for each participant, $e_{i}=\Pr(X_{i}=1\mid Z_{i})$, is often estimated from the fitted model $\mbox{logit}(e_{i})=G_{i}^{T}\alpha$, where $G_{i}=g(Z_{i})$ is a vector of predictors for participant $i$ for some user-specified function $g$ of $Z_{i}$ and $\alpha$ is the vector of regression coefficients from the propensity model.
+The predicted propensity score for each participant is calculated as $\widehat{e}_{i}=e(Z_{i},\widehat{\alpha})=\mbox{logit}^{-1}(G_{i}^{T}\hat{\alpha})$ where $\hat{\alpha}$ is the maximum likelihood estimate (MLE) of $\alpha$.
 The IPTW is then estimated for each participant as $\hat{W}_{i}=X_{i}\hat{e}_{i}^{-1}+(1-X_{i})(1-\hat{e}_{i})^{-1}$.
@@
 That is, $a_{x}(Z_{i},\hat{\gamma})=\hat{E}(Y_{i}\mid Z_{i},X_{i}=x)$ for $x\in\{0,1\}$, where $\hat{\gamma}$ are the MLEs of the parameters from the assumed outcome model.
-Then, the classic AIPW estimator for the $ACE$ is
+Then, $ACE$ is estimated by
 $\widehat{DR}_{C}=\hat{\mu}_{C}^{1}-\hat{\mu}_{C}^{0}$ (1)
@@
 2.2.2 Weighted regression AIPW estimator
-An alternative AIPW estimator incorporates the estimated propensity scores when fitting the outcome regression model.
-The parameters of the $Y\mid Z,X$ model are estimated with IPTW-weighted maximum likelihood estimation.
+For the weighted regression AIPW estimator, the parameters of the $Y\mid Z,X$ model are estimated with IPTW-weighted maximum likelihood estimation.
 Pseudo-outcomes are obtained for each participant under both exposures, i.e., $b_{x}(Z_{i},\hat{\beta})=\hat{E}(Y_{i}\mid Z_{i},X_{i}=x)$ for $x\in\{0,1\}$, where $\hat{\beta}$ are the MLEs from the weighted outcome model.
@@
 where $\hat{\mu}_{WR}^{x}=n^{-1}\sum_{i=1}^{n}b_{x}(Z_{i},\hat{\beta})$ for $x\in\{0,1\}$.
-The weighted regression AIPW estimator (2) has been evaluated previously (Kang and Schafer,, 2007; Robins et al.,, 2007; Vansteelandt and Keiding,, 2011; Gabriel et al.,, 2023) and is expected to be more stable than (1) when IPTWs are extreme due to bounding within the parameter space (Kang and Schafer,, 2007; Robins et al.,, 2007; Vansteelandt et al.,, 2012).
+The estimator (2) has been evaluated previously and is expected to be more stable than (1) when IPTWs are extreme due to bounding within the parameter space (Kang and Schafer,, 2007; Robins et al.,, 2007; Vansteelandt et al.,, 2012).
 2.2.3 TMLE
-TMLE is an alternative doubly robust estimator that consists of a two-stage process.
+TMLE consists of a two-stage process.
 In the first stage, $Y\mid Z,X$ is modeled as in Section 2.2.1, and $a_{x}(Z_{i},\hat{\gamma})$ are calculated for $x\in\{0,1\}$.
-In the second stage, referred to as the targeting stage (van der Laan and Rubin,, 2006), corrections to outcome regression predictions from stage one are made that incorporate the estimated propensity scores.
+In the second stage, referred to as the targeting stage, corrections to outcome regression predictions from stage one are made that incorporate the estimated propensity scores.
 Different methods have been proposed for fitting the targeting model; here, the weighted regression approach is considered, as it is thought to handle random positivity violations better (van der Laan et al.,, 2011).
-Specifically, the models $\mbox{logit}(Y_{i})=\eta_{x}+\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}$ are fit for $x\in\{0,1\}$.
-Parameters $\eta_{x}$ are estimated using weighted maximum-likelihood, with weights $(1-X)\hat{W}_{i}$ and $X\hat{W}_{i}$ for the $x=0$ and $x=1$ targeting model, respectively.
-In the targeting models, logit-transformed pseudo-outcomes ($\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}$) for each participant are included as offset terms, and MLEs $\hat{\eta}_{0}$ and $\hat{\eta}_{1}$ can be thought of as corrections to these pseudo-outcomes (Zivich et al., 022b, ).
+First, assuming $Y$ is binary, the models $\mbox{logit}\{\Pr(Y_{i}=1\mid Z_{i},X_{i}=x)\}=\eta_{x}+\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}$ are fit for $x\in\{0,1\}$.
+Note the only parameter in this model is $\eta_{x}$, as $\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}$ is a fixed offset.
+Parameters $\eta_{x}$ are estimated using weighted maximum likelihood, with weights $(1-X)\hat{W}_{i}$ and $X\hat{W}_{i}$ for the $x=0$ and $x=1$ models, respectively.
 The stage two pseudo-outcomes are computed as $c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})=\mbox{expit}[\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}+\hat{\eta}_{x}]$.
-The TMLE estimator of $ACE$ is
+Then, the TMLE estimator of $ACE$ is
 $\widehat{DR}_{TMLE}=\hat{\mu}_{TMLE}^{1}-\hat{\mu}_{TMLE}^{0}$ (3)
 where $\hat{\mu}_{TMLE}^{x}=n^{-1}\sum_{i=1}^{n}c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})$ for $x\in\{0,1\}$.
-Note that for continuous outcomes, such as birth weight considered in Section 3, the outcome is scaled prior to implementing TMLE, i.e., $Y_{i}^{*}=(Y_{i}-a)/(b-a)$ is defined for each participant, where $(a,b)$ are the bounds of $Y$.
-Then, $Y_{i}$ is replaced with $Y_{i}^{*}$ in the steps above, and $c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})=\mbox{expit}[\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}+\hat{\eta}_{x}](b-a)+a$ re-scales the outcome back to the original bounds of $Y$.
-TMLE methods were originally proposed by van der Laan and Rubin, (2006) and have been extended and applied in a number of settings (Stitelman et al.,, 2012; Balzer et al.,, 2023; van der Laan et al.,, 2018; van der Laan and Gruber,, 2012).
+For continuous outcomes, a logit model is still recommended for the targeting stage to ensure bounding within the parameter space (Gruber and van der Laan,, 2010; Gruber and Van Der Laan,, 2012).
+Here, the outcome is scaled prior to implementing TMLE, i.e., $Y_{i}^{*}=(Y_{i}-a)/(b-a)$ is defined, where $(a,b)$ are the bounds of $Y$.
+Then, $Y_{i}$ is replaced with $Y_{i}^{*}$ in the steps above, i.e., $E(Y_{i}^{*}\mid Z_{i},X_{i}=x)$ is modeled with a logit link, and $c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})=\mbox{expit}[\mbox{logit}\{a_{x}(Z_{i},\hat{\gamma})\}+\hat{\eta}_{x}](b-a)+a$ re-scales the outcome back to the original bounds of $Y$.
+TMLE methods were developed by van der Laan and Rubin, (2006) and are frequently paired with machine learning techniques for nuisance modeling.
 When parametric models are used in both stages of estimation, (3) is expected to be similar to (2) (Tran et al.,, 2019).
 2.3 Variance Estimation
-Variance estimation for doubly robust estimators (1), (2), and (3) is not straightforward, because the variance of each estimator depends on estimating the parameters of the nuisance models.
-Here, we consider two methods for variance estimation: the commonly used influence function based variance estimator and the empirical sandwich variance estimator.
-Both variance estimators can be used to construct Wald-typed CIs for the $ACE$.
+Variance estimation for doubly robust estimators (1), (2), and (3) is not straightforward, because the variance of each estimator depends on estimating the parameters of the working models.
+Here, three methods for variance estimation are considered, each of which can be used to construct Wald-typed CIs for the $ACE$.
 2.3.1 Influence function based variance estimator
 Estimators (1), (2), and (3) are suggested by the efficient influence function (Luque-Fernandez et al.,, 2018; Gabriel et al.,, 2023).
-The commonly used influence function based variance estimator has the form (Glynn and Quinn,, 2010; Gruber and Van Der Laan,, 2012)
-$\hat{V}(\widehat{DR})=n^{-2}\sum_{i=1}^{n}\hat{I}_{i}^{2}$ (4)
-where $\hat{I}_{i}=\hat{e}_{i}^{-1}X_{i}Y_{i}-(1-\hat{e}_{i})^{-1}(1-X_{i})Y_{i}-\{\hat{e}_{i}^{-1}(1-\hat{e}_{i})^{-1}(X_{i}-\hat{e}_{i})\}\{(1-\hat{e}_{i})\hat{Y}_{i}^{1}+\hat{e}_{i}\hat{Y}_{i}^{0}\}-\widehat{DR}$.
+The commonly used influence function based variance estimator has the form $\hat{V}(\widehat{DR})_{IF}=n^{-2}\sum_{i=1}^{n}\hat{I}_{i}^{2}$, where $\hat{I}_{i}=\hat{e}_{i}^{-1}X_{i}Y_{i}-(1-\hat{e}_{i})^{-1}(1-X_{i})Y_{i}-\{\hat{e}_{i}^{-1}(1-\hat{e}_{i})^{-1}(X_{i}-\hat{e}_{i})\}\{(1-\hat{e}_{i})\hat{Y}_{i}^{1}+\hat{e}_{i}\hat{Y}_{i}^{0}\}-\widehat{DR}$ (Gruber and Van Der Laan,, 2012).
 Here, pseudo-outcomes $\hat{Y}_{i}^{x}$ equal $a_{x}(Z_{i},\hat{\gamma})$, $b_{x}(Z_{i},\hat{\beta})$, and $c_{x}(O_{i},\hat{\gamma},\hat{\eta}_{x})$ for the classic AIPW, weighted regression AIPW, and TMLE methods, respectively, and $\widehat{DR}$ equals (1), (2), and (3) for the three corresponding estimators.
-The estimators (4) are consistent for the variance of (1), (2), and (3) when both propensity score and outcome models are correctly specified (Daniel,, 2014; Gruber and Van Der Laan,, 2012).
-When one of the two nuisance models is misspecified, the influence function based variance estimator is not consistent.
-As noted in the introduction, the estimator (4) is conservative when the weight model is correctly specified but the outcome model is misspecified (Muñoz and Van Der Laan,, 2012), and may be conservative or anti-conservative when the outcome model is correctly specified but the weight model is misspecified (Laan and Robins,, 2003; Muñoz and Van Der Laan,, 2012).
-In empirical studies, the variance estimator (4) for the classic AIPW estimator (1) was found to be unstable for small sample sizes ($n<1000$), even under correct model specification (Funk et al.,, 2011).
-Similar instability has been reported for (4) with TMLE ($n=100$) (van der Laan and Gruber,, 2011).
+The influence function based variance estimators $\hat{V}(\widehat{DR})_{IF}$ are consistent for the variance of (1), (2), and (3) when both propensity score and outcome models are correctly specified (Daniel,, 2014; Gruber and Van Der Laan,, 2012).
+When one of the two working models is misspecified, the influence function based variance estimator is not consistent.
+As noted in the introduction, the estimator $\hat{V}(\widehat{DR})_{IF}$ is conservative when the weight model is correctly specified but the outcome model is misspecified, and may be conservative or anti-conservative when the outcome model is correctly specified but the weight model is misspecified.
 2.3.2 Empirical sandwich variance estimator
-An alternative to the influence function based variance estimator (4) is the empirical sandwich variance estimator.
 The estimators (1), (2), and (3) can each be expressed as the solution to a set of unbiased estimating equations.
 The M-estimator for each set of estimating equations, $\hat{\theta}$, is the solution (for $\theta$) to $\sum_{i=1}^{n}\psi_{q}(O_{i},\theta)=0$, where $\psi_{q}(O_{i},\theta)$ are the estimating functions corresponding to each of the three estimators $q\in\{1,2,3\}$ with parameter vector $\theta$.
-Parameter vectors and estimating functions for (1), (2), and (3) can be specified as follows.
+Parameter vectors and estimating functions for each estimator are provided in Appendix A.
+[…]
+2.3.3 Nonparametric Bootstrap
+The nonparametric bootstrap can also be used to estimate the variance of (1), (2), and (3).
+While variations exist, the nonparametric bootstrap typically involves selecting $B$ independent, with replacement resamples of size $n$ from the observed sample.
+The estimator of interest is applied to each resample $b$ to obtain $\widehat{DR}_{b}$, where $\widehat{DR}_{b}$ equals (1), (2), or (3) applied to resample $b$.
+Then, $\hat{V}(\widehat{DR})_{NB}=\sum_{b=1}^{B}\{\widehat{DR}_{b}-\widehat{DR}^{*}\}^{2}/(B-1)$, where $\widehat{DR}^{*}=B^{-1}\sum_{b=1}^{B}\widehat{DR}_{b}$.
+For large samples, the nonparametric bootstrap is expected to provide valid estimates of the variance for smooth functions of solutions to smooth estimating equations (Davison,, 1997), including (1), (2), and (3).
+Therefore, the nonparametric bootstrap is a doubly robust variance estimator.
+However, the bootstrap’s performance may suffer in certain settings, e.g., in the presence of outliers (Davison,, 1997).
+3 Example: Improving Pregnancy Outcomes with Progesterone
+The doubly robust estimators described in Section 2.2 were applied to IPOP data for estimation of the effect of maternal anemia on birth weight.
+The exposure, maternal anemia, was defined as having a baseline hemoglobin level below 10.5 g/dL.
+The IPOP data were limited to the 782 participants (98%) with no prior preterm births and with measured birth weights and baseline hemoglobin values.
+Note 14 participants experienced stillbirth.
+Birth weights for these participants were included in the analysis; we return to this issue below.
+It was assumed that potential outcomes are independent of the exposure conditional on the covariates listed in Appendix Table A1.
+To demonstrate differences in the variance estimators presented in Section 2.3, each doubly robust estimator was applied with three sets of model specifications.
+First, the covariates in Appendix Table A1 were included in both propensity and outcome models.
+Then, a naive outcome model was fit that included only the exposure variable, while the propensity model included the full set of covariates.
+Under this specification, the outcome model was reasonably thought to be misspecified due to its simplicity.
+Finally, a naive propensity model was fit that included only an intercept term, i.e., assuming the propensity score is constant, while the outcome model included the full set of covariates.
+This propensity model was similarly thought to be misspecified.
+For all estimators, maternal height was modeled using restricted cubic splines with four knots placed at the 5th, 35th, 65th, and 95th percentiles.
+The empirical sandwich variance estimator was computed using geex in R (Saul and Hudgens,, 2020) and delicatessen in Python (Zivich et al., 022c, ).
+The bootstrap was based on 5000 resamples.
+Point estimates and corresponding 95% CIs for each method and model specification are presented in Table 1, with 95% CI half-widths compared in Figure 1.
+The doubly robust methods provided similar point estimates of the $ACE$ for all model specifications, with estimates of approximately -40g, though all 95% CIs included zero.
+Precision estimates based on the empirical sandwich variance estimator and the influence function based variance estimator were similar when the full covariate set was included in both working models.
+Under the naive outcome model, standard errors for the influence function based variance estimator were larger than those for the empirical sandwich variance estimator.
+When the propensity model was naive, standard errors for the influence function based variance estimator were smaller than those for the empirical sandwich variance estimator.
+Note there was more fluctuation in 95% CI half-widths across model specifications for the influence function based variance estimator than the empirical sandwich variance estimator (Figure 1).
+Bootstrap standard errors were larger than those of the empirical sandwich and influence function based variance estimators.
+This may be due to the presence of extreme birth weights in the data (Appendix Figure A1).
+There are notable limitations associated with this analysis.
+First, the $ACE$ is likely not the most informative estimand to address this research question.
+The $ACE$ contrasts average birth weight under settings where all women have anemia and no women have anemia.
+It is challenging to envision a setting where all women would be susceptible to anemia, so the estimators (1), (2), and (3) may lack a meaningful causal interpretation.
+While these statistical quantities might still be beneficial, more informative estimands could also be considered to address this research question, including a comparison of $E(Y^{0})$ with the natural course, $E(Y)$ (Hubbard and Van der Laan,, 2008) or contrasts of stochastic intervention distributions (Kennedy,, 2019).
+Additionally, as with all observational studies, there may be uncontrolled confounding not captured in the observed set of covariates.
+This concern may be heightened in settings where causal consistency is also questionable (Hernán and VanderWeele,, 2011), i.e., there may be multiple interventions to modify anemia status.
+Finally, the analysis included participants who experienced stillbirth.
+Future work could reexamine this problem using alternative approaches for competing events.
+Table 1: Estimated effect of maternal anemia on birth weight by model specification and estimator
+$\widehat{DR}$ ES-SE ES-95% CI IF-SE IF-95% CI NB-SE NB-95% CI
+Full Covariate Set
+Classic AIPW -37 56 (-147, 73) 58 (-151, 76) 63 (-161, 86)
+Wtd regression AIPW -41 56 (-151, 69) 56 (-151, 68) 61 (-160, 78)
+TMLE -37 56 (-147, 73) 58 (-151, 76) 63 (-160, 85)
+Naive Outcome
+Classic AIPW -36 57 (-148, 76) 61 (-156, 84) 64 (-162, 90)
+Wtd regression AIPW -36 57 (-148, 77) 61 (-156, 84) 64 (-161, 90)
+TMLE -36 57 (-148, 77) 61 (-156, 84) 63 (-160, 89)
+Naive Propensity
+Classic AIPW -41 58 (-154, 73) 56 (-150, 69) 58 (-155, 74)
+Wtd regression AIPW -47 57 (-159, 65) 54 (-153, 59) 59 (-162, 68)
+TMLE -41 58 (-154, 73) 56 (-150, 69) 59 (-155, 74)
+  • Note: SE=standard error, ES=empirical sandwich, IF=influence function, NB=nonparametric bootstrap, CI=confidence interval, Wtd=weighted.
+The naive outcome model included only the exposure, and the naive propensity model included only an intercept.
+The nonparametric bootstrap was based on 5000 resamples.
+Figure 1: 95% confidence interval (CI) half-widths by model specification and estimator. FC=Full Covariate Set, NO=Naive Outcome Model, NP=Naive Propensity Model
+4 Simulation Study
+A simulation study was conducted to compare the empirical properties of the influence function based variance estimator, the empirical sandwich variance estimator, and the nonparametric bootstrap in conjunction with each of the three doubly robust methods discussed in Section 2.
+Simulations were conducted with $n=800$, similar to the sample size in the example in Section 3.
+4.1 Simulation Setup
+Three covariates ($Z_{1}$, $Z_{2}$, $Z_{3}$), the exposure ($X$), and potential outcomes ($Y^{0}$ and $Y^{1}$) were simulated.
+The covariate $Z_{1}$ was distributed Normal with mean $155$ and standard deviation $7.6$.
+Two binary covariates $Z_{2}$ and $Z_{3}$ were simulated from Bernoulli distributions with means $0.25$ and $0.75$, respectively.
+The exposure $X$ was simulated from a Bernoulli distribution with mean $\mbox{expit}(15-0.1Z_{1}+2.5Z_{2}-Z_{3}-0.02Z_{1}Z_{2}+0.005Z_{1}Z_{3})$.
+Potential outcomes $Y^{1}$ and $Y^{0}$ under exposure and no exposure, respectively, were simulated from a normal distribution with mean $E(Y^{x})=1000+11.5Z_{1}+100Z_{2}-15Z_{1}Z_{2}+25x-5.5xZ_{1}-30xZ_{2}+5xZ_{1}Z_{2}$ and standard deviation $\sigma=400$ for $x\in\{0,1\}$.
+Under this data generating mechanism, the $ACE$ was approximately $-60$.
+The marginal distributions of the exposure and outcome were modeled after the example in Section 3.
+To examine performance of the estimators under the null, simulations were also conducted where $E(Y^{1})$ was equal to $E(Y^{0})$, as defined above, such that $ACE=0$.
+The estimators $\widehat{DR}_{C}$, $\widehat{DR}_{WR}$, and $\widehat{DR}_{TMLE}$ were applied to 5000 simulated samples.
+Estimators were computed four ways: (1) propensity and outcome models correctly specified, (2) propensity models correctly specified but outcome models misspecied, (3) outcome models correctly specified but propensity models misspecified, and (4) both models misspecified.
+Misspecified propensity models included only an intercept and a linear term for $(Z_{1}-155)^{2}$, and misspecified outcome models included only an intercept and linear terms for $X$ and $(Z_{1}-155)^{2}$.
+Each of the three variance estimators described in Section 2.3 was applied.
+For the nonparametric bootstrap, 1000 resamples were included in each iteration for estimation of the bootstrap standard error, excluding any resamples where working models failed to converge.
+Simulation results were summarized by empirical bias, average standard error (ASE), empirical standard error (ESE), standard error ratio (SER=ASE/ESE), and empirical 95% CI coverage.
+The ratio of the variance estimate for each simulation and the ESE, i.e., the variance ratio, was also summarized.
+That is, $VR=se_{s}/ESE$, where $se_{s}=\sqrt{\hat{V}(\widehat{DR})}$ for a given doubly robust estimator, model specification, and variance estimator for simulation $s$.
+4.2 Simulation Results
+Both point and variance estimators performed as expected in simulations.
+The results of the simulation study are presented in Table 2 and Figures 2 and 3.
+When at least one model was correctly specified, the classic AIPW, weighted regression AIPW, and TMLE displayed minimal bias, but all were substantially biased when both working models were misspecified.
+Under correct specification of both working models, all three variance estimators tracked closely with the ESE, resulting in SERs close to one and CIs attaining the nominal level of coverage.
+When both models were misspecified, all three variance estimators tracked with the ESE, but bias was substantial, resulting in CIs with below nominal coverage.
+Differences between variance estimators are apparent for scenarios where only one working model was correctly specified.
+When either the outcome model or the propensity model was misspecified, CIs based on the empirical sandwich variance estimator and the nonparametric bootstrap attained the nominal level of coverage, and estimated standard errors generally tracked closely with the ESE.
+In Figures 2 and 3, note variance ratios for the empirical sandwich estimator and the nonparametric bootstrap are clustered around one for these scenarios.
+In contrast, the influence function based variance estimator was empirically biased when either working model was misspecified.
+As expected, it tended to overestimate the variance when the outcome model was misspecified, leading to SERs above one and conservative CI coverage.
+When the propensity model was misspecified but the outcome model was correctly specified, the influence function based variance estimator underestimated the variance, resulting in SERs below one.
+Bias in the influence function based variance estimator was more pronounced under the null.
+Under outcome model misspecification, SERs exceeded 2.0, resulting in approximately 100% CI coverage.
+Under propensity misspecification, SERs were 0.89, resulting in CIs with below nominal coverage.
+The empirical sandwich variance and nonparametric bootstrap estimators both demonstrated the expected doubly robust variance property, but there was a notable difference in the performance of the estimators.
+The nonparametric bootstrap generally demonstrated more variation than the empirical sandwich estimator, as evidenced by increased spread and more extreme outliers in Figures 2 and 3.
+Additional simulations were conducted with $\sigma\in\{200,600\}$, with a binary outcome, and with $n=2000$.
+Twelve scenarios were considered.
+Details and results of the additional scenarios are provided in Appendix C.
+In summary, the simulation study demonstrated the theoretical properties explained in Section 2.
+That is, the empirical sandwich variance estimator and the nonparametric bootstrap were empirically unbiased when at least one of the two working models was correctly specified.
+The influence function based variance estimator is not consistent under misspecification of either working model.
+The influence function based variance estimator was conservative when the outcome model was misspecified and was generally anti-conservative when the propensity model was misspecified.
+The magnitude and direction of bias under misspecified working models varied across simulation scenarios.
+Figure 2: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, continuous outcome, n=800, $\sigma=400$, 5000 simulations.
+$ACE$ was approximately -60.
+Black squares denote the mean variance ratio (=SER).
+Results exclude one simulation where models failed to converge.
+The $0.33\%$ of correct model specification simulations, $4.02\%$ of misspecified outcome model simulations, and $0.004\%$ of misspecified propensity model simulations where the ratio was above 1.2 or below 0.8 are not displayed.
+Figure 3: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, continuous outcome, n=800, $\sigma=400$, 5000 simulations under the null.
+Black squares denote the mean variance ratio (=SER).
+The $0.01\%$ of correct model specification simulations and $1.72\%$ of misspecified outcome model simulations where the ratio was above 2.75 or below 0.5 are not displayed.
+Table 2: Simulation summary results, continuous outcome, $n=800$, $\sigma=400$, $5000$ simulations. Bias, ESE, SER, and 95% CI coverage calculated for the ACE.
+Scenario Estimator Bias ESE SER, ES Cov, ES (%) SER, NB Cov, NB (%) SER, IF Cov, IF (%)
+$ACE\approx-60$
+CS Classic 0.4 58.4 0.99 95 1.00 95 1.00 95
+WR 0.4 58.3 0.99 95 1.00 95 0.99 95
+TMLE 0.4 58.3 0.99 95 1.00 95 1.00 95
+MO Classic -0.4 60.0 0.99 95 1.02 96 1.07 97
+WR -1.6 59.2 0.99 95 1.00 95 1.05 96
+TMLE -0.5 59.7 0.99 95 1.01 95 1.06 96
+MP Classic 0.3 57.8 1.00 95 1.00 95 0.97 94
+WR 0.3 57.8 1.00 95 1.00 95 0.97 94
+TMLE 0.3 57.8 1.00 95 1.00 95 0.97 94
+MB Classic -23.8 57.0 1.00 92 1.00 92 1.00 92
+WR -23.8 57.0 1.00 92 1.00 92 1.00 92
+TMLE -23.8 57.0 1.00 92 1.00 92 1.00 92
+$ACE=0$
+CS Classic 0.1 35.0 0.96 94 1.00 95 0.97 94
+WR 0.1 34.8 0.96 94 0.98 94 0.95 94
+TMLE 0.1 35.0 0.96 94 0.99 95 0.97 94
+MO Classic 1.4 48.0 0.93 94 1.07 96 2.10 100
+WR 2.5 44.8 0.94 93 1.02 95 2.18 100
+TMLE 1.6 46.7 0.93 94 1.04 96 2.12 100
+MP Classic 0.0 33.7 0.98 95 0.99 95 0.89 92
+WR -0.1 33.7 0.98 95 0.99 95 0.89 92
+TMLE 0.0 33.7 0.98 95 0.99 95 0.89 92
+MB Classic 158.8 75.9 0.99 44 0.99 44 0.99 44
+WR 158.8 75.9 0.99 44 0.99 44 0.99 44
+TMLE 158.8 75.9 0.99 44 0.99 44 0.99 44
+  • ESE=empirical standard error; SER=standard error ratio (ASE/ESE), where ASE=average estimated standard error; Cov = 95% confidence interval coverage, i.e., the proportion of simulated samples for which the 95% CI included ACE; ES=empirical sandwich variance estimator; NB=nonparametric bootstrap variance estimator, IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MP=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
+Results exclude 1 simulation where models did not converge.
+5 Discussion
+Doubly robust estimators have gained popularity due to their ability to provide consistent point estimates when either an outcome or propensity model is correctly specified.
+Here, the commonly-used influence function based variance estimator is compared with the empirical sandwich variance estimator and the nonparametric bootstrap in conjunction with three doubly robust estimators: the classic AIPW estimator, the weighted regression AIPW estimator, and TMLE.
+For estimation of the average causal effect with observational data, the influence function based variance estimator is consistent only when both outcome and propensity models are correctly specified.
+In contrast, both the empirical sandwich variance estimator and the nonparametric bootstrap are doubly robust variance estimators.
+As such, confidence intervals constructed from these estimators are expected to provide nominal CI coverage when either model is correctly specified.
+This paper only considers variance estimation of the average causal effect with observational data.
+The influence function based variance estimator can be consistent under misspecification of one working model in some settings.
+For example, this variance estimator is consistent under outcome model misspecification in the randomized setting under additional assumptions (Williams et al.,, 2022; Van Lancker et al.,, 2024) and for estimation of some non-degenerate estimands (Haneuse and Rotnitzky,, 2013).
+While here consideration was limited to finite-dimensional parametric modeling approaches, machine learning approaches are commonly applied with doubly robust estimators.
+These approaches allow for more complex functional forms for continuous covariates than fully parametric approaches and include higher order interaction terms that may not be included in investigator-specified parametric models, often leading to estimators that are more robust to model misspecification (Zivich et al.,, 2022).
+However, convergence of machine learning algorithms is typically slower and consistent variance estimation is more challenging than with parametric modeling approaches.
+Machine learning methods are typically not compatible with the estimating equations approach discussed in this paper, though alternative methods have been developed for doubly robust variance estimation in this context for TMLE estimators (Benkeser et al.,, 2017).
+Use of the empirical sandwich variance estimator in conjunction with doubly robust estimators is not new, though its doubly robust property has not been emphasized.
+Lunceford and Davidian, (2004) discuss use of the empirical sandwich variance estimator for estimating the variance of the classic AIPW estimator and note that it tends to be more stable than the influence function based variance estimator.
+However, based on the systematic review discussed in the introduction (Smith et al.,, 2023), the empirical sandwich variance estimator does not appear to be widely used with TMLE.
+While some existing software packages compute empirical sandwich variance estimators for doubly robust estimators (e.g., the “causaltrt” procedure in SAS and the “dr” procedure in Stata), other popular software packages compute the influence function based variance estimator but not the empirical sandwich variance estimator (e.g., the AIPW and tmle packages in R and the zEpid package in Python).
+We hope this work will allow for easier implementation of doubly robust point and variance estimation.
+Supplementary Material
+R and Python code for computing the different estimators along with the corresponding standard error estimators is available at https://github.com/bonnieshook/DoublyRobustVariance.
+Acknowledgements
+This work was supported by the National Institutes of Health under award numbers R01 AI157758, K01 AI182506, R01 AI085073, P30 AI050410, and K01 AI177102.
+The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
+The authors thank Drs.
+Bradley Saul and Michael Hudgens at the University of North Carolina for coding support and helpful suggestions, respectively.
+Appendix A: parameter vectors and estimating functions for three doubly robust estimators with the M-estimation approach
 Classic AIPW estimator
@@
 $\psi_{2}(O_{i};\theta)=\begin{bmatrix}\psi_{\alpha}\\ \psi_{\beta}\\ \psi_{\mu_{WR1}}\\ \psi_{\mu_{WR0}}\\ \psi_{WR}\end{bmatrix}=\begin{bmatrix}\{X_{i}-\mbox{expit}(G_{i}^{T}\alpha)\}G_{i}\\ W_{i}\{Y_{i}-\phi^{-1}(H_{i}^{T}\beta)\}H_{i}\\ b_{1}(Z_{i},\beta)-\mu_{WR}^{1}\\ b_{0}(Z_{i},\beta)-\mu_{WR}^{0}\\ \mu_{WR}^{1}-\mu_{WR}^{0}-DR_{WR}\end{bmatrix}$
-where $\psi_{\beta}$ is the vector of IPTW-weighted score functions for the outcome model.
-The estimating functions $\psi_{\mu_{WR1}}$ and $\psi_{\mu_{WR0}}$ are for the causal means $\mu_{WR}^{1}$ and $\mu_{WR}^{0}$, respectively, and $\psi_{WR}$ is a delta method transformation of the causal means for estimation of the $ACE$ using the weighted AIPW estimator $\widehat{DR}_{WR}$.
-Gabriel et al., (2023) demonstrate unbiasedness of this set of estimating equations when at least one of the nuisance models is correctly specified.
+where $\psi_{\beta}$ is the vector of IPTW-weighted score functions for the outcome model and $b_{x}(Z_{i},\beta)=\phi^{-1}(H_{i}(x,Z_{i}))^{\top}\beta$ for $x\in\{0,1\}$.
+The estimating functions $\psi_{\mu_{WR1}}$ and $\psi_{\mu_{WR0}}$ are for the causal means $\mu_{WR}^{1}$ and $\mu_{WR}^{0}$, respectively, and $\psi_{WR}$ is a delta method transformation of the causal means for estimation of the $ACE$ using the weighted regression AIPW estimator $\widehat{DR}_{WR}$.
+Gabriel et al., (2023) demonstrate unbiasedness of this set of estimating equations when at least one of the working models is correctly specified.
 TMLE
-With TMLE, $\theta=[\alpha,\gamma,\eta_{1},\eta_{0},\mu_{TMLE}^{1},\mu_{TMLE}^{0},DR_{TMLE}]$. For continuous outcomes, the estimating function $\psi_{3}$ is specified as
+For TMLE, $\theta=[\alpha,\gamma,\eta_{1},\eta_{0},\mu_{TMLE}^{1},\mu_{TMLE}^{0},DR_{TMLE}]$. For continuous outcomes, the estimating function $\psi_{3}$ is specified as
 $\psi_{3}(O_{i};\theta)=\begin{bmatrix}\psi_{\alpha}\\ \psi_{\gamma}\\ \psi_{t1}\\ \psi_{t0}\\ \psi_{\mu_{TMLE1}}\\ \psi_{\mu_{TMLE0}}\\ \psi_{TMLE}\end{bmatrix}=\begin{bmatrix}\{X_{i}-\mbox{expit}(G_{i}^{T}\alpha)\}G_{i}\\ \{Y^{*}_{i}-\phi^{-1}(H_{i}^{T}\gamma)\}H_{i}\\ X_{i}e_{i}^{-1}(Y^{*}_{i}-\mbox{expit}[\eta_{1}+\mbox{logit}\{a_{1}(Z_{i},\gamma)\}])\\ (1-X_{i})(1-e_{i})^{-1}(Y^{*}_{i}-\mbox{expit}[\eta_{0}+\mbox{logit}\{a_{0}(Z_{i},\gamma)\}])\\ c_{1}(O_{i},\gamma,\eta_{1})-\mu_{TMLE}^{1}\\ c_{0}(O_{i},\gamma,\eta_{0})-\mu_{TMLE}^{0}\\ \mu_{TMLE}^{1}-\mu_{TMLE}^{0}-DR_{TMLE}\end{bmatrix}$
@@
 The estimating functions $\psi_{t1}$ and $\psi_{t0}$ are for the two targeting models, while $\psi_{\mu_{TMLE1}}$ and $\psi_{\mu_{TMLE0}}$ are for the causal means $\mu_{TMLE}^{1}$ and $\mu_{TMLE}^{0}$, respectively.
-Note $c_{x}(O_{i},\gamma,\eta_{x})=\mbox{expit}[\mbox{logit}\{a_{x}(Z_{i},\gamma)\}+\eta_{x}](b-a)+a$ are the pseudo-outcomes following the targeting step, as outlined in Section 2.2.3.
+Note $c_{x}(O_{i},\gamma,\eta_{x})=\mbox{expit}[\mbox{logit}\{a_{x}(Z_{i},\gamma)\}+\eta_{x}](b-a)+a$ are the pseudo-outcomes following the targeting step, as outlined in Section 2.2.3 of the main text.
 Finally, $\psi_{TMLE}$ is a delta method transformation of the causal means for estimation of the $ACE$ using $\widehat{DR}_{TMLE}$.
-Unbiasedness of $\psi_{3}(O_{i};\theta)$ when at least one nuisance model is correctly specified is demonstrated in the Appendix.
-[…]
-3 Example: Improving Pregnancy Outcomes with Progesterone
-Prior research has shown links between maternal anemia and low birth weight (Azizah et al.,, 2022; Figueiredo et al.,, 2018).
-The relationship between anemia and birth weight is particularly important to understand for women living with HIV, as they experience anemia more frequently than women without HIV (Levine et al.,, 2001) and are more prone to adverse birth outcomes (Tukei et al.,, 2021).
-The effect of maternal anemia on birth weight among women living with HIV was estimated with data from the IPOP trial.
-The IPOP trial was a randomized, double-blind, placebo-controlled trial conducted in Lusaka, Zambia.
-Participants were pregnant women aged 18 or older living with HIV on (or intending to initiate) antiretroviral therapy (ART) who had viable singleton pregnancies at less than 24 weeks of gestation at baseline.
-The 800 enrolled participants were randomized with equal probability to either a placebo or 17 alpha-hydroxyprogesterone caproate (17P) treatment to evaluate the efficacy of 17P for prevention of preterm birth.
-Additional details about the IPOP trial have been published previously (Price et al.,, 2021).
-While IPOP was designed to examine the effect of 17P on preterm birth or stillbirth, covariates collected during the trial allow for examination of other causal questions, assuming that the identification assumptions hold.
-The three doubly robust estimators described in Section 2.2 were applied to IPOP data for estimation of the $ACE$ of maternal anemia on birth weight.
-The exposure, maternal anemia, was defined as having a baseline hemoglobin level below 10.5 g/dL.
-The IPOP data were limited to the 782 participants (98%) with no prior preterm births and with measured birth weights and baseline hemoglobin values.
-Note 14 participants experienced stillbirth.
-Birth weights for these participants were included in the analysis; we return to this issue in the discussion.
-It was assumed that potential outcomes are independent of the exposure conditional on the following baseline covariates: trimester (1 or 2), maternal age group (18-24, 25-29, 30-34, 35-39, 40+), number of prior births (0, 1, 2, 3, 4+), alcohol use during pregnancy (yes or no), ART use (yes or no), and maternal height (cm) (Table 1).
-At baseline, no IPOP participants reported tobacco or drug use during pregnancy, so these covariates were not included.
-Primary analyses were conducted using R and were independently replicated in Python.
-As discussed in the introduction, application of doubly robust estimators for estimating the $ACE$ of maternal anemia on birth weight may offer precision advantages over IPTW estimators, but the correct model specifications are unknown.
-To demonstrate differences in the influence function based variance estimator and the empirical sandwich variance estimator presented in Section 2.3, each of the doubly robust estimators were applied with three sets of model specifications.
-First, the covariates in Table 1 were included in both propensity and outcome models.
-Then, a naive outcome model was fit that included only the exposure variable (anemia status), while the propensity model included the full set of covariates.
-Under this specification, the outcome model was reasonably thought to be misspecified due to its simplicity.
-Finally, a naive propensity model was fit that included only an intercept term, while the outcome model included the full set of covariates.
-Due to the simplicity of the propensity model, it was reasonably thought to be misspecified.
-For all estimators, maternal height was modeled using restricted cubic splines with four knots placed at the 5th, 35th, 65th, and 95th percentiles (Perperoglou et al.,, 2019).
-The empirical sandwich variance estimator was computed using geex in R (Saul and Hudgens,, 2020) and delicatessen in Python (Zivich et al., 022c, ).
-Confidence intervals for the $ACE$ were constructed based on the influence function based variance estimator and the empirical sandwich variance estimator.
-Estimates for each of the three methods and model specifications along with corresponding 95% CIs are presented in Table 2, with 95% CI half-widths compared in Figure 1.
-IPOP participants with and without anemia at baseline had similar self-reported alcohol use during pregnancy and similar age and height distributions.
-Participants without anemia were more likely to have one or more prior births, be in their first trimester, and be on ART.
-The three doubly robust methods provided similar estimates of the $ACE$ for all model specifications examined, with average birth weights if all women had anemia were approximately 40g lower than birth weights if no women had anemia, though all 95% CIs included zero.
-Precision estimates based on the empirical sandwich variance estimator and the influence function based variance estimator were similar when the full covariate set was included in both the outcome and propensity models.
-Under the naive outcome model, standard errors for the influence function based variance estimator were larger than those for the empirical sandwich variance estimator, resulting in wider CIs.
-When the propensity model was naive, standard errors for the influence function based variance estimator were smaller than those for the empirical sandwich variance estimator, resulting in narrower CIs.
-Note there was more fluctuation in 95% CI half-widths across model specifications for the influence function based variance estimator than the empirical sandwich variance estimator (Figure 1).
-Table 1: Baseline characteristics of IPOP participants by anemia status
-No anemia $n=664$ Anemia $n=118$
-Age category 18-24 156 (23%) 28 (24%)
-25-29 190 (29%) 33 (28%)
-30-34 192 (29%) 31 (26%)
-35-39 100 (15%) 21 (18%)
-40+ 26 (4%) 5 (4%)
-Number of prior births 0 125 (19%) 30 (25%)
-1 160 (24%) 25 (21%)
-2 157 (24%) 28 (24%)
-3 125 (19%) 18 (15%)
-4+ 97 (15%) 17 (14%)
-First trimester 120 (18%) 12 (10%)
-ART use 648 (98%) 109 (92%)
-Alcohol use during pregnancy 64 (10%) 9 (8%)
-Height Median (Q1,Q3) 155 (150, 160) 154 (150, 159)
-Mean (SD) 155 (8) 153 (8)
-Min, Max 131, 185 132, 170
-  • Note: Anemia defined as baseline hemoglobin below 10.5 g/dL
-Table 2: Estimated effect of maternal anemia on birth weight by model specification and estimator
-ACE ES-SE ES-95% CI IF-SE IF-95% CI
-Full Covariate Set
-Classic AIPW -37 56 (-147, 73) 58 (-151, 76)
-Weighted regression AIPW -41 56 (-151, 69) 56 (-151, 68)
-TMLE -37 56 (-147, 73) 58 (-151, 76)
-Naive Outcome Model
-Classic AIPW -36 57 (-148, 76) 61 (-156, 84)
-Weighted regression AIPW -36 57 (-148, 77) 61 (-156, 84)
-TMLE -36 57 (-148, 77) 61 (-156, 84)
-Naive Propensity Model
-Classic AIPW -41 58 (-154, 73) 56 (-150, 69)
-Weighted regression AIPW -47 57 (-159, 65) 54 (-153, 59)
-TMLE -40 58 (-154, 73) 56 (-150, 69)
-  • Note: SE=standard error, ES=empirical sandwich, IF=influence function, CI=confidence interval.
-The naive outcome model included only the exposure, and the naive propensity model included only an intercept.
-Figure 1: 95% confidence interval (CI) half-widths by model specification and estimator. FC=Full Covariate Set, NO=Naive Outcome Model, NP=Naive Propensity Model
-4 Simulation Study
-A simulation study was conducted to compare the empirical properties of the influence function based variance estimator and the empirical sandwich variance estimator in conjunction with each of the three doubly robust methods discussed in Section 2.
-4.1 Simulation Setup
-Simulations were designed based on the example in Section 3, which aimed to estimate the effect of maternal anemia on birth weight among women with HIV.
-The primary simulations were conducted with $n=800$ participants, similar to the sample size in the example.
-Results for a sample size of $n=2000$ are included in the Supporting Information.
-The simulation study was conducted in R with code provided on GitHub.
-Three covariates ($Z_{1}$, $Z_{2}$, $Z_{3}$), the exposure ($X$), and potential outcomes ($Y^{0}$ and $Y^{1}$) were simulated.
-The covariate $Z_{1}$ was distributed Normal with mean $155$ and standard deviation $7.6$.
-Two binary covariates $Z_{2}$ and $Z_{3}$ were simulated from Bernoulli distributions with means $0.25$ and $0.75$, respectively.
-The exposure $X$ was simulated from a Bernoulli distribution with mean $\mbox{expit}(15-0.1Z_{1}+2.5Z_{2}-1Z_{3}-0.02Z_{1}Z_{2}+0.005Z_{1}Z_{3})$.
-Potential outcomes $Y^{0}$ and $Y^{1}$ under exposure and no exposure, respectively, were simulated from a normal distribution with mean $E(Y^{x})=1000+11.5Z_{1}+100Z_{2}-15Z_{1}Z_{2}+25x-5.5xZ_{1}-30xZ_{2}+20xZ_{1}Z_{2}$ and standard deviation $\sigma=400$ for $x\in\{0,1\}$.
-Under this data generating mechanism, the $ACE$ was approximately $-60$.
-Additional simulations were conducted with $\sigma=200$ with results in the Supporting Information.
-The estimators $\widehat{DR}_{C}$, $\widehat{DR}_{WR}$, and $\widehat{DR}_{TMLE}$ were applied to 5000 simulated samples to estimate the $ACE$.
-Estimators were computed four ways: (1) with both propensity and outcome models correctly specified, (2) with propensity models correctly specified but outcome models misspecied, (3) with outcome models correctly specified but propensity models misspecified, and (4) with both models misspecified.
-Misspecified propensity models included only an intercept and a linear term for $(Z_{1}-155)^{2}$, and misspecified outcome models included only an intercept and linear terms for $X$ and $(Z_{1}-155)^{2}$.
-[…]
-4.2 Simulation Results
-Both point and variance estimators performed as expected in simulations.
-The results of the simulation study for $n=800$ and $\sigma=400$ are presented in Table 3 and Figure 2.
-When at least one model was correctly specified, the classic AIPW, weighted regression AIPW, and TMLE displayed minimal bias, but all were substantially biased when both nuisance models were misspecified.
-Note bias was further reduced under correct specification of at least one model for $n=2000$ (Table 4).
-Under correct specification of both nuisance models, the empirical sandwich and influence function based variance estimators tracked closely with the ESE, resulting in SERs close to one and CIs attaining the nominal level of coverage.
-When both models were misspecified, both variance estimators tracked with the ESE, but bias was substantial, resulting in CIs with below nominal coverage.
-Differences between the variance estimators are apparent when examining the simulation study results for scenarios where one nuisance model was correctly specified and the other was misspecified.
-When either the outcome model or the propensity model was misspecified, CIs based on the empirical sandwich variance estimator attained the nominal level of coverage, and estimated standard errors tracked closely with the ESE.
-In Figure 2, note variance ratios for the empirical sandwich estimator are closely clustered around one for these scenarios.
-In contrast, the influence function based variance estimator was empirically biased when either nuisance model was misspecified.
-As expected, it tended to overestimate the variance when the outcome model was misspecified, leading to estimated standard errors that tracked above the ESE, SERs above one, and conservative CIs that exceeded the nominal level of coverage.
-When the propensity model was misspecified but the outcome model was correctly specified, the influence function based variance estimator underestimated the variance, resulting in estimated standard errors below the ESE and SERs below one.
-The biases observed for the influence function based variance estimator were also present when $n=2000$ (Table 4 and Figure 3).
-In the $\sigma=200$ scenarios, the empirical sandwich variance estimator performance was similar to the $\sigma=400$ scenario, but the performance of the influence function based variance estimator varied (Tables 5-6 and Figures 4-5).
-Specifically, when the propensity model was misspecified but the outcome model was correctly specified, the influence function based variance estimator demonstrated minimal empirical bias, resulting in SERs close to one.
-In summary, the simulation study demonstrated the theoretical properties explained in Section 2.
-That is, the empirical sandwich variance estimator was consistent when at least one of the two nuisance models was correctly specified, and performed well in simulations.
-The influence function based variance estimator is not consistent under misspecification of either nuisance model.
-In line with expectations, the influence function based variance estimator was conservative when the outcome model was misspecified.
-For this data generating mechanism, it was anti-conservative when the propensity model was misspecified for $\sigma=400$ and exhibited minimal bias when $\sigma=200$.
-Figure 2: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, n=800, $\sigma=400$, 5000 simulations.
-Black squares denote the mean variance ratio (=SER).
-Results exclude one simulation where models failed to converge.
-Note the $0.3\%$ of correct model specification simulations and $4.5\%$ of misspecified outcome model simulations where the ratio was above 1.2 are not displayed.
-Table 3: Simulation summary results, $n=800$, $\sigma=400$, $5000$ simulations. Bias, Relative Bias, ASE, ESE, SER, and 95% CI coverage calculated for the ACE.
-Scenario Estimator Bias Relative Bias (%) ESE ASE, ES SER, ES Cov, ES (%) ASE, IF SER, IF Cov, IF (%)
-CS Classic 0.4 0.6 58.4 58.0 0.99 95 58.1 1.00 95
-WR 0.4 0.6 58.3 57.7 0.99 95 57.6 0.99 95
-TMLE 0.4 0.6 58.3 57.9 0.99 95 58.1 1.00 95
-MO Classic -0.4 0.6 60.0 59.2 0.99 95 64.5 1.07 97
-WR -1.7 2.8 59.2 58.4 0.99 95 62.4 1.05 96
-TMLE -0.5 0.8 59.7 58.9 0.99 95 63.0 1.06 96
-MW Classic 0.3 0.5 57.8 57.6 1.00 95 55.8 0.97 94
-WR 0.2 0.4 57.7 57.6 1.00 95 55.8 0.97 94
-TMLE 0.3 0.5 57.8 57.6 1.00 95 55.8 0.97 94
-MB Classic -23.8 39.8 57.0 56.8 1.00 92 56.8 1.00 92
-WR -23.8 39.8 57.0 56.8 1.00 92 56.8 1.00 92
-TMLE -23.8 39.8 57.0 56.8 1.00 92 56.8 1.00 92
-  • ASE=average estimated standard error; ESE=empirical standard error; SER=standard error ratio (ASE/ESE); Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MW=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
-Results exclude one simulation where models failed to converge.
-5 Discussion
-Doubly robust estimators have increased in popularity in the field of causal inference due to their ability to provide consistent point estimates when either an outcome or propensity model is correctly specified (Smith et al.,, 2023).
-In this paper, the commonly-used influence function based variance estimator is compared with the empirical sandwich variance estimator in conjunction with three doubly robust estimators: the classic AIPW estimator, the weighted regression AIPW estimator, and TMLE.
-These three estimators can be specified as solutions to sets of unbiased estimating equations, and thus their variances can be consistently estimated by the empirical sandwich variance estimator.
-Because the estimating equations remain unbiased when either the outcome or propensity model is correctly specified, the empirical sandwich variance estimator is a doubly robust variance estimator.
-In contrast, the influence function based variance estimator is consistent only when both outcome and propensity models are correctly specified.
-This discrepancy may be of particular importance in settings where the functional form of one of the nuisance models is known but the other is not.
-For example, with randomized trials, randomization to treatment means that the exposure mechanism is known.
-Doubly robust estimators can offer improved efficiency in this setting, but some of that efficiency gain may not be realized with application of the influence function based variance estimator, which is conservative if the outcome model is misspecified.
-While consideration in this paper was limited to finite-dimensional parametric modeling approaches, machine learning approaches can also be used with doubly robust estimators (Kennedy et al.,, 2016; van der Laan and Gruber,, 2011, 2012; van der Laan et al.,, 2018; Balzer et al.,, 2023).
-Machine learning approaches allow for more complex functional forms for continuous covariates than fully parametric approaches and include higher order interaction terms that may not be included in investigator-specified parametric models, often leading to estimators that are more robust to model misspecification (Zivich et al.,, 2022).
-However, convergence of machine learning algorithms is typically slower and consistent variance estimation is more challenging than with parametric modeling approaches.
-Machine learning methods are typically not compatible with the estimating equations approach discussed in this paper, though alternative methods have been developed for doubly robust variance estimation in this context for TMLE estimators (Benkeser et al.,, 2017).
-In the example, the effect of anemia on birth weight among women with HIV was estimated by applying the three doubly robust methods to data from the IPOP randomized trial.
-Average birth weights if all women had anemia were estimated to be approximately 40g lower than birth weights if no women had anemia, though all 95% CIs included zero.
-The analysis included 14 participants who experienced stillbirth.
-Because stillbirth is a competing outcome for live birth weight, future work could reexamine this problem using alternative approaches for competing events.
-Use of the empirical sandwich variance estimator in conjunction with doubly robust estimators is not new, though its doubly robust property has not been emphasized.
-Lunceford and Davidian, (2004) discuss use of the empirical sandwich variance estimator for estimating the variance of the classic AIPW estimator and note that it tends to be more stable than the influence function based variance estimator.
-However, based on the systematic review discussed in the introduction (Smith et al.,, 2023), the influence function based variance estimator is frequently applied and the empirical sandwich variance estimator does not appear to be widely used with TMLE.
-While some existing software packages compute empirical sandwich variance estimators for doubly robust estimators (e.g., the “causaltrt” procedure in SAS and the “dr” procedure in Stata), other popular software packages compute the influence function based variance estimator but not the empirical sandwich variance estimator (e.g., the AIPW and tmle packages in R and the zEpid package in Python).
-We provide example R and Python code for estimating the variance of each of the three doubly robust estimators with the empirical sandwich variance estimator, which we hope will allow for easier implementation of doubly robust point and variance estimation.
-Supplementary Material
-R and Python code for computing the different estimators along with the corresponding standard error estimators is available at https://github.com/bonnieshook/DoublyRobustVariance.
-Acknowledgements
-This work was supported by US National Institutes of Health under award numbers R01 AI157758, R01 AI085073, P30 AI050410, and K01 AI177102.
-The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
-The authors thank Bradley Saul at the University of North Carolina at Chapel Hill for coding support.
-Appendix
-Consistency proof of TMLE estimator under correct specification of either the outcome or propensity model
+Unbiasedness of $\psi_{3}(O_{i};\theta)$ when at least one working model is correctly specified is demonstrated in Appendix B.
+Appendix B: Consistency proof of TMLE estimator under correct specification of either the outcome or propensity model
 Recall the TMLE estimating function $\psi_{3}$ is specified as
 $\psi_{3}(O_{i};\theta)=\begin{bmatrix}\psi_{\alpha}\\ \psi_{\gamma}\\ \psi_{t1}\\ \psi_{t0}\\ \psi_{\mu_{TMLE1}}\\ \psi_{\mu_{TMLE0}}\\ \psi_{TMLE}\end{bmatrix}=\begin{bmatrix}\{X_{i}-\mbox{expit}(G_{i}^{T}\alpha)\}G_{i}\\ \{Y^{*}_{i}-\phi^{-1}(H_{i}^{T}\gamma)\}H_{i}\\ X_{i}e_{i}^{-1}(Y^{*}_{i}-\mbox{expit}[\eta_{1}+\mbox{logit}\{a_{1}(Z_{i},\gamma)\}])\\ (1-X_{i})(1-e_{i})^{-1}(Y^{*}_{i}-\mbox{expit}[\eta_{0}+\mbox{logit}\{a_{0}(Z_{i},\gamma)\}])\\ c_{1}(O_{i},\gamma,\eta_{1})-\mu_{TMLE}^{1}\\ c_{0}(O_{i},\gamma,\eta_{0})-\mu_{TMLE}^{0}\\ \mu_{TMLE}^{1}-\mu_{TMLE}^{0}-DR_{TMLE}\end{bmatrix}$
-where $\theta=[\alpha,\gamma,\eta_{1},\eta_{0},\mu_{TMLE}^{1},\mu_{TMLE}^{0},DR_{TMLE}]$ is the vector of parameters.
-Let
+where $\theta=[\alpha,\gamma,\eta_{1},\eta_{0},\mu_{TMLE}^{1},\mu_{TMLE}^{0},DR_{TMLE}]$ is the vector of parameters. Let
 $\tilde{\theta}=[\tilde{\alpha},\tilde{\gamma},\tilde{\eta}_{1},\tilde{\eta}_{0},\tilde{\mu}_{TMLE}^{1},\tilde{\mu}_{TMLE}^{0},\widetilde{DR}_{TMLE}]$ be the solution of $E\{\psi_{3}(O_{i};\theta)\}=0$ and $\widehat{DR}_{TMLE}$ be the proposed TMLE estimator of ACE, obtained by solving the estimating equation $n^{-1}\sum_{i}\psi_{3}(O_{i};\theta)=0$.
@@
 In conclusion, if either the outcome model or the propensity score model is correctly specified, $\widetilde{DR}_{TMLE}=\tilde{\mu}_{TMLE}^{1}-\tilde{\mu}_{TMLE}^{0}=\mu^{1}-\mu^{0}=ACE$ which proves the stated result.
-Table 4: Simulation summary results, $n=2000$, $\sigma=400$, $5000$ simulations. Bias, Relative Bias, ASE, ESE, SER, and 95% CI coverage calculated for the ACE.
-Scenario Estimator Bias Relative Bias (%) ESE ASE, ES SER, ES Cov, ES (%) ASE, IF SER, IF Cov, IF (%)
-CS Classic -0.1 0.2 36.6 36.7 1.00 95 36.8 1.00 95
-WR -0.1 0.2 36.6 36.6 1.00 95 36.6 1.00 95
-TMLE -0.1 0.2 36.6 36.7 1.00 95 36.8 1.00 95
-MO Classic -0.3 0.6 37.2 37.4 1.01 95 40.6 1.09 97
-WR -0.9 1.6 37.0 37.1 1.00 95 39.6 1.07 96
-TMLE -0.4 0.6 37.1 37.3 1.01 95 39.8 1.07 96
-MW Classic -0.2 0.3 36.2 36.4 1.01 95 35.3 0.98 94
-WR -0.2 0.3 36.2 36.4 1.01 95 35.3 0.98 95
-TMLE -0.2 0.3 36.2 36.4 1.01 95 35.3 0.98 94
-MB Classic -23.8 39.8 35.9 35.9 1.00 90 35.9 1.00 90
-WR -23.8 39.8 35.9 35.9 1.00 90 35.9 1.00 90
-TMLE -23.8 39.8 35.9 35.9 1.00 90 35.9 1.00 90
-  • ASE=average estimated standard error; ESE=empirical standard error; SER=standard error ratio (ASE/ESE); Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MW=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
-Table 5: Simulation summary results, $n=800$, $\sigma=200$, $5000$ simulations. Bias, Relative Bias, ASE, ESE, SER, and 95% CI coverage calculated for the ACE.
-Scenario Estimator Bias Relative Bias (%) ESE ASE, ES SER, ES Cov, ES (%) ASE, IF SER, IF Cov, IF (%)
-CS Classic 0.3 0.6 50.2 50.0 1.00 95 50.0 1.00 95
-WR 0.3 0.5 50.1 49.9 1.00 95 49.9 1.00 95
-TMLE 0.3 0.5 50.2 50.0 1.00 95 50.0 1.00 95
-MO Classic -0.4 0.6 52.0 51.5 0.99 95 57.2 1.10 97
-WR -1.6 2.7 51.3 50.7 0.99 95 55.1 1.07 96
-TMLE -0.5 0.8 51.6 51.1 0.99 95 55.6 1.08 96
-MW Classic 0.3 0.5 50.0 49.9 1.00 95 49.4 0.99 95
-WR 0.3 0.4 50.0 49.9 1.00 95 49.4 0.99 95
-TMLE 0.3 0.5 50.0 49.9 1.00 95 49.4 0.99 95
-MB Classic -23.6 39.4 50.7 50.4 0.99 92 50.4 0.99 92
-WR -23.6 39.4 50.7 50.4 0.99 92 50.4 0.99 92
-TMLE -23.5 39.3 50.7 50.4 0.99 92 50.4 0.99 92
-  • ASE=average estimated standard error; ESE=empirical standard error; SER=standard error ratio (ASE/ESE); Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MW=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
-Results exclude 14 simulations where models did not converge.
-Table 6: Simulation summary results, $n=2000$, $\sigma=200$, $5000$ simulations. Bias, Relative Bias, ASE, ESE, SER, and 95% CI coverage calculated for the ACE.
-Scenario Estimator Bias Relative Bias (%) ESE ASE, ES SER, ES Cov, ES (%) ASE, IF SER, IF Cov, IF (%)
-CS Classic 0.0 0.0 31.6 31.6 1.00 95 31.7 1.00 95
-WR 0.0 0.1 31.5 31.6 1.00 95 31.6 1.00 95
-TMLE 0.0 0.0 31.6 31.6 1.00 95 31.7 1.00 95
-MO Classic -0.2 0.4 32.3 32.4 1.00 95 36.0 1.11 97
-WR -0.8 1.4 32.1 32.1 1.00 95 34.9 1.09 97
-TMLE -0.3 0.4 32.1 32.3 1.00 95 35.1 1.09 97
-MW Classic 0.0 0.0 31.4 31.6 1.00 95 31.2 0.99 95
-WR 0.0 0.0 31.4 31.6 1.00 95 31.2 0.99 95
-TMLE 0.0 0.0 31.4 31.6 1.00 95 31.2 0.99 95
-MB Classic -23.6 39.5 32.0 31.9 1.00 89 31.9 1.00 89
-WR -23.6 39.5 32.0 31.9 1.00 89 31.9 1.00 89
-TMLE -23.6 39.5 32.0 31.9 1.00 89 31.9 1.00 89
-  • ASE=average estimated standard error; ESE=empirical standard error; SER=standard error ratio (ASE/ESE); Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MW=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
-Figure 3: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, n=2000, $\sigma=400$, 5000 simulations.
+Appendix C: Additional Simulation Scenarios
+Additional Simulation Setup
+In addition to simulations described in Section 4 of the main text, 10 additional simulation scenarios were considered:
+  1. 1.
+    continuous outcome, $n=800$, $\sigma=200$, $ACE\approx-60$
+  2. 2.
+    continuous outcome, $n=800$, $\sigma=600$, $ACE\approx-60$
+  3. 3.
+    continuous outcome, $n=2000$, $\sigma=200$, $ACE\approx-60$
+  4. 4.
+    continuous outcome, $n=2000$, $\sigma=400$, $ACE\approx-60$
+  5. 5.
+    continuous outcome, $n=2000$, $\sigma=600$, $ACE\approx-60$
+  6. 6.
+    continuous outcome, $n=2000$, $\sigma=400$, $ACE=0$ (null)
+  7. 7.
+    binary outcome, $n=800$, $ACE\approx 0.045$
+  8. 8.
+    binary outcome, $n=800$, $ACE=0$ (null)
+  9. 9.
+    binary outcome, $n=2000$, $ACE\approx 0.045$
+  10. 10.
+    binary outcome, $n=2000$, $ACE=0$ (null)
+The continuous outcome scenarios followed the same data generating mechanism described in Section 4 of the main text, with the specified sample size and value of $\sigma$.
+For binary outcome scenarios, the distributions of $Z_{1}$, $Z_{2}$, and $Z_{3}$ were the same as the continuous outcome scenarios, but the exposure and outcomes differed.
+The exposure $X$ was simulated from a Bernoulli distribution with mean $\mbox{expit}(16-0.1Z_{1}+2.5Z_{2}-Z_{3}-0.02Z_{1}Z_{2}+0.005Z_{1}Z_{3})$.
+Potential outcomes $Y^{0}$ and $Y^{1}$ under no exposure and exposure, respectively, were simulated from Bernoulli distributions with means $\mbox{expit}\{-2+0.5(Z_{1}-155)+0.75Z_{2}-0.8Z_{2}(Z_{1}-155)-0.2(Z_{1}-155)(1-x)-0.1Z_{2}(1-x)+Z_{2}(Z_{1}-155)(1-x)\}$ for $x\in\{0,1\}$.
+Under this data generating mechanism, $ACE\approx 0.045$.
+To examine performance of the estimators under the null, simulations were also conducted for the binary outcome where $E(Y^{1})$ was equal to $E(Y^{0})$, as defined above, such that $ACE=0$.
+Additional Simulation Results
+Results of the additional simulation scenarios are presented in Appendix Tables A2-A11 and Appendix Figures A2-A11.
+When $n=2000$, empirical bias in point estimates, empirical sandwich variance estimates, and nonparametric bootstrap estimates under correct specification of at least one model was further reduced.
+However, an increase in the sample size did not mitigate bias in the influence function based variance estimator when one model was misspecified.
+For continuous outcomes, under misspecification of the outcome model, the conservative property of the influence function based variance estimator decreased as $\sigma$ increased.
+Under misspecification of the propensity model, influence function based variance estimators became increasingly anti-conservative as $\sigma$ increased.
+For the binary outcome, the conservative property of the influence function based variance estimator under outcome model misspecification was more pronounced than for the continuous outcome.
+Notably, under misspecified propensity models, the influence function based variance estimator was slightly conservative when there was non-null effect but slightly anti-conservative under the null.
+These findings highlight that the direction and magnitude of bias under propensity model misspecification vary across data generating mechanisms.
+Table A1: Baseline characteristics of IPOP participants by anemia status
+No anemia $n=664$ Anemia $n=118$
+Age category 18-24 156 (23%) 28 (24%)
+25-29 190 (29%) 33 (28%)
+30-34 192 (29%) 31 (26%)
+35-39 100 (15%) 21 (18%)
+40+ 26 (4%) 5 (4%)
+Number of prior births 0 125 (19%) 30 (25%)
+1 160 (24%) 25 (21%)
+2 157 (24%) 28 (24%)
+3 125 (19%) 18 (15%)
+4+ 97 (15%) 17 (14%)
+First trimester 120 (18%) 12 (10%)
+ART use 648 (98%) 109 (92%)
+Alcohol use during pregnancy 64 (10%) 9 (8%)
+Height (cm) Median (Q1,Q3) 155 (150, 160) 154 (150, 159)
+Mean (SD) 155 (8) 153 (8)
+Min, Max 131, 185 132, 170
+  • Note: Anemia defined as baseline hemoglobin below 10.5 g/dL
+Table A2: Simulation summary results, continuous outcome, $n=800$, $\sigma=200$, $5000$ simulations. Bias, ESE, SER, and 95% CI coverage calculated for the ACE.
+Scenario Estimator Bias ESE SER, ES Cov, ES (%) SER, NB Cov, NB (%) SER, IF Cov, IF (%)
+CS Classic 0.4 50.1 1.00 95 1.00 95 1.00 95
+WR 0.4 50.1 1.00 95 1.00 95 1.00 95
+TMLE 0.3 50.1 1.00 95 1.00 95 1.00 95
+MO Classic -0.3 51.9 0.99 95 1.02 96 1.10 97
+WR -1.6 51.3 0.99 95 1.00 95 1.07 96
+TMLE -0.4 51.6 0.99 95 1.01 95 1.08 96
+MP Classic 0.4 50.0 1.00 95 1.00 95 0.99 95
+WR 0.3 50.0 1.00 95 1.00 95 0.99 95
+TMLE 0.3 50.0 1.00 95 1.00 95 0.99 95
+MB Classic -23.5 50.7 0.99 92 1.00 92 0.99 92
+WR -23.5 50.7 0.99 92 1.00 92 0.99 92
+TMLE -23.5 50.7 0.99 92 1.00 92 0.99 92
+  • ESE=empirical standard error; SER=standard error ratio (ASE/ESE), where ASE=average estimated standard error; Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; NB=nonparametric bootstrap variance estimator, IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MP=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
+Results exclude 15 simulations where models failed to converge.
+Table A3: Simulation summary results, continuous outcome, $n=800$, $\sigma=600$, $5000$ simulations. Bias, ESE, SER, and 95% CI coverage calculated for the ACE.
+Scenario Estimator Bias ESE SER, ES Cov, ES (%) SER, NB Cov, NB (%) SER, IF Cov, IF (%)
+CS Classic 0.4 70.1 0.99 94 1.00 95 0.99 95
+WR 0.4 69.9 0.98 94 1.00 95 0.98 94
+TMLE 0.4 70.1 0.99 94 1.00 95 0.99 95
+MO Classic -0.4 71.5 0.98 95 1.01 95 1.05 96
+WR -1.7 70.6 0.98 95 1.00 95 1.03 96
+TMLE -0.5 71.2 0.98 95 1.00 95 1.04 95
+MP Classic 0.3 68.9 0.99 95 1.00 95 0.95 93
+WR 0.2 68.9 0.99 95 1.00 95 0.95 93
+TMLE 0.3 68.9 0.99 95 1.00 95 0.95 93
+MB Classic -24.1 66.2 1.00 93 1.00 93 1.00 93
+WR -24.1 66.2 1.00 93 1.00 93 1.00 93
+TMLE -24.1 66.2 1.00 93 1.00 93 1.00 93
+  • ESE=empirical standard error; SER=standard error ratio (ASE/ESE), where ASE=average estimated standard error; Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; NB=nonparametric bootstrap variance estimator, IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MP=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
+Table A4: Simulation summary results, continuous outcome, $n=2000$, $\sigma=200$, $5000$ simulations. Bias, ESE, SER, and 95% CI coverage calculated for the ACE.
+Scenario Estimator Bias ESE SER, ES Cov, ES (%) SER, NB Cov, NB (%) SER, IF Cov, IF (%)
+CS Classic 0.0 31.6 1.00 95 1.00 95 1.00 95
+WR 0.0 31.5 1.00 95 1.00 95 1.00 95
+TMLE 0.0 31.6 1.00 95 1.00 95 1.00 95
+MO Classic -0.2 32.3 1.00 95 1.01 95 1.11 97
+WR -0.8 32.1 1.00 95 1.01 95 1.09 97
+TMLE -0.2 32.1 1.00 95 1.01 95 1.09 97
+MP Classic 0.0 31.4 1.00 95 1.00 95 0.99 95
+WR 0.0 31.4 1.00 95 1.00 95 0.99 95
+TMLE 0.0 31.4 1.00 95 1.00 95 0.99 95
+MB Classic -23.6 32.0 1.00 89 1.00 89 1.00 89
+WR -23.6 32.0 1.00 89 1.00 89 1.00 89
+TMLE -23.6 32.0 1.00 89 1.00 89 1.00 89
+  • ESE=empirical standard error; SER=standard error ratio (ASE/ESE), where ASE=average estimated standard error; Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; NB=nonparametric bootstrap variance estimator, IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MP=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
+Table A5: Simulation summary results, continuous outcome, $n=2000$, $\sigma=400$, $5000$ simulations. Bias, ESE, SER, and 95% CI coverage calculated for the ACE.
+Scenario Estimator Bias ESE SER, ES Cov, ES (%) SER, NB Cov, NB (%) SER, IF Cov, IF (%)
+CS Classic -0.1 36.6 1.00 95 1.01 95 1.00 95
+WR -0.1 36.6 1.00 95 1.00 95 1.00 95
+TMLE -0.1 36.6 1.00 95 1.01 95 1.00 95
+MO Classic -0.3 37.2 1.01 95 1.01 95 1.09 97
+WR -0.9 37.0 1.00 95 1.01 95 1.07 96
+TMLE -0.4 37.1 1.01 95 1.01 95 1.07 96
+MP Classic -0.2 36.2 1.01 95 1.01 95 0.98 94
+WR -0.2 36.2 1.01 95 1.01 95 0.98 95
+TMLE -0.2 36.2 1.01 95 1.01 95 0.98 94
+MB Classic -23.8 35.9 1.00 90 1.00 90 1.00 90
+WR -23.8 35.9 1.00 90 1.00 90 1.00 90
+TMLE -23.8 35.9 1.00 90 1.00 90 1.00 90
+  • ESE=empirical standard error; SER=standard error ratio (ASE/ESE), where ASE=average estimated standard error; Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; NB=nonparametric bootstrap variance estimator, IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MP=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
+Table A6: Simulation summary results, continuous outcome, $n=2000$, $\sigma=600$, $5000$ simulations. Bias, ESE, SER, and 95% CI coverage calculated for the ACE.
+Scenario Estimator Bias ESE SER, ES Cov, ES (%) SER, NB Cov, NB (%) SER, IF Cov, IF (%)
+CS Classic -0.2 43.8 1.00 95 1.01 95 1.00 95
+WR -0.2 43.8 1.00 95 1.00 95 1.00 95
+TMLE -0.2 43.8 1.00 95 1.01 95 1.00 95
+MO Classic -0.5 44.3 1.01 95 1.01 96 1.07 96
+WR -1.1 44.0 1.00 95 1.01 95 1.05 96
+TMLE -0.5 44.2 1.00 95 1.01 95 1.05 96
+MP Classic -0.4 43.1 1.01 95 1.01 95 0.96 94
+WR -0.4 43.1 1.01 95 1.01 95 0.96 94
+TMLE -0.4 43.1 1.01 95 1.01 95 0.96 94
+MB Classic -24.0 41.6 1.00 91 1.00 91 1.00 91
+WR -24.0 41.6 1.00 91 1.00 91 1.00 91
+TMLE -24.0 41.6 1.00 91 1.00 91 1.00 91
+  • ESE=empirical standard error; SER=standard error ratio (ASE/ESE), where ASE=average estimated standard error; Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; NB=nonparametric bootstrap variance estimator, IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MP=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
+Table A7: Simulation summary results, continuous outcome, $n=2000$, $\sigma=400$, $5000$ simulations under the null. Bias, ESE, SER, and 95% CI coverage calculated for the ACE.
+Scenario Estimator Bias ESE SER, ES Cov, ES (%) SER, NB Cov, NB (%) SER, IF Cov, IF (%)
+CS Classic -0.2 21.6 1.00 95 1.00 95 1.00 95
+WR -0.2 21.6 0.99 95 1.00 95 0.99 95
+TMLE -0.2 21.6 0.99 95 1.00 95 1.00 95
+MO Classic 0.3 28.9 0.96 95 1.01 96 2.20 100
+WR 0.9 27.3 0.97 94 1.00 95 2.26 100
+TMLE 0.4 28.3 0.97 95 1.01 95 2.21 100
+MP Classic -0.3 21.0 1.00 95 1.01 95 0.91 93
+WR -0.3 21.0 1.00 95 1.01 95 0.91 93
+TMLE -0.3 21.0 1.00 95 1.00 95 0.91 93
+MB Classic 159.4 47.5 1.00 8 1.00 9 1.00 8
+WR 159.4 47.5 1.00 8 1.00 9 1.00 8
+TMLE 159.4 47.5 1.00 8 1.00 9 1.00 8
+  • ESE=empirical standard error; SER=standard error ratio (ASE/ESE), where ASE=average estimated standard error; Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; NB=nonparametric bootstrap variance estimator, IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MP=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was 00; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
+Table A8: Simulation summary results, binary outcome, $n=800$, $5000$ simulations. Bias, ESE, SER, and 95% CI coverage calculated for the ACE.
+Scenario Estimator Bias ESE SER, ES Cov, ES (%) SER, NB Cov, NB (%) SER, IF Cov, IF (%)
+CS Classic 0.0 2.6 0.99 95 0.99 95 0.99 95
+WR 0.0 2.6 0.99 95 0.99 95 0.99 95
+TMLE 0.0 2.6 0.99 95 0.99 95 0.99 95
+MO Classic 0.0 2.9 0.97 94 1.00 95 1.28 99
+WR -0.1 2.8 0.97 94 1.07 95 1.24 99
+TMLE 0.0 2.9 0.97 94 0.99 95 1.22 98
+MP Classic 0.0 2.6 0.99 95 0.99 95 1.02 95
+WR 0.0 2.6 0.99 95 0.99 95 1.02 95
+TMLE 0.0 2.6 0.99 95 0.99 95 1.02 95
+MB Classic -15.7 3.2 0.99 0 0.99 0 0.99 0
+WR -15.7 3.2 0.99 0 0.99 0 0.99 0
+TMLE -15.7 3.2 0.99 0 0.99 0 0.99 0
+  • ESE=empirical standard error; SER=standard error ratio (ASE/ESE), where ASE=average estimated standard error; Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; NB=nonparametric bootstrap variance estimator, IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MP=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $0.045$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
+Results exclude eight simulations where models failed to converge.
+Table A9: Simulation summary results, binary outcome, $n=800$, $5000$ simulations under the null. Bias, ESE, SER, and 95% CI coverage calculated for the ACE.
+Scenario Estimator Bias ESE SER, ES Cov, ES (%) SER, NB Cov, NB (%) SER, IF Cov, IF (%)
+CS Classic 0.0 2.4 0.99 95 0.99 95 0.99 95
+WR 0.0 2.4 0.99 95 0.99 95 0.99 95
+TMLE 0.0 2.4 0.99 95 1.06 95 0.99 95
+MO Classic 0.0 2.5 1.00 96 1.03 96 1.49 100
+WR -0.2 2.5 1.00 95 1.04 96 1.39 99
+TMLE 0.0 2.5 1.00 95 1.02 96 1.39 100
+MP Classic 0.0 2.4 0.99 95 1.00 95 0.98 94
+WR 0.0 2.4 0.99 95 1.00 95 0.98 95
+TMLE 0.0 2.4 0.99 95 1.06 95 0.98 95
+MB Classic -20.1 3.1 0.99 0 0.99 0 0.99 0
+WR -20.1 3.1 0.99 0 0.99 0 0.99 0
+TMLE -20.1 3.1 0.99 0 0.99 0 0.99 0
+  • ESE=empirical standard error; SER=standard error ratio (ASE/ESE), where ASE=average estimated standard error; Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; NB=nonparametric bootstrap variance estimator, IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MP=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was 00; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
+Results exclude eight simulations where models failed to converge.
+Table A10: Simulation summary results, binary outcome, $n=2000$, $5000$ simulations. Bias, ESE, SER, and 95% CI coverage calculated for the ACE.
+Scenario Estimator Bias ESE SER, ES Cov, ES (%) SER, NB Cov, NB (%) SER, IF Cov, IF (%)
+CS Classic 0.0 1.6 0.99 95 1.00 95 0.99 95
+WR 0.0 1.6 0.99 95 1.00 95 0.99 95
+TMLE 0.0 1.6 0.99 95 1.00 95 0.99 95
+MO Classic 0.0 1.8 1.00 95 1.01 95 1.30 99
+WR -0.1 1.7 0.99 95 1.10 95 1.26 99
+TMLE 0.0 1.8 0.99 95 1.00 95 1.25 99
+MP Classic 0.0 1.6 0.99 95 0.99 95 1.03 96
+WR 0.0 1.6 0.99 95 0.99 95 1.03 96
+TMLE 0.0 1.6 0.99 95 0.99 95 1.03 96
+MB Classic -15.6 2.0 1.00 0 1.00 0 1.00 0
+WR -15.6 2.0 1.00 0 1.00 0 1.00 0
+TMLE -15.6 2.0 1.00 0 1.00 0 1.00 0
+  • ESE=empirical standard error; SER=standard error ratio (ASE/ESE), where ASE=average estimated standard error; Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; NB=nonparametric bootstrap variance estimator, IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MP=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $0.045$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
+Results exclude four simulations where models failed to converge.
+Table A11: Simulation summary results, binary outcome, $n=2000$, $5000$ simulations under the null. Bias, ESE, SER, and 95% CI coverage calculated for the ACE.
+Scenario Estimator Bias ESE SER, ES Cov, ES (%) SER, NB Cov, NB (%) SER, IF Cov, IF (%)
+CS Classic 0.0 1.5 0.98 94 0.98 94 0.98 94
+WR 0.0 1.5 0.98 94 0.98 95 0.98 94
+TMLE 0.0 1.5 0.98 94 0.98 95 0.98 94
+MO Classic 0.0 1.6 0.99 95 1.00 95 1.49 100
+WR -0.1 1.6 0.99 95 1.03 95 1.39 99
+TMLE 0.0 1.6 0.99 95 1.00 95 1.39 99
+MP Classic 0.0 1.5 0.98 95 0.98 95 0.97 94
+WR 0.0 1.5 0.98 95 0.98 95 0.97 94
+TMLE 0.0 1.5 0.98 95 0.98 95 0.97 94
+MB Classic -20.0 2.0 0.98 0 0.98 0 0.98 0
+WR -20.0 2.0 0.98 0 0.98 0 0.98 0
+TMLE -20.0 2.0 0.98 0 0.98 0 0.98 0
+  • ESE=empirical standard error; SER=standard error ratio (ASE/ESE), where ASE=average estimated standard error; Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; NB=nonparametric bootstrap variance estimator, IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MP=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was 00; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
+Results exclude one simulation where models failed to converge.
+Figure A1: Boxplot of birth weights among $n=782$ IPOP participants by anemia status.
+Figure A2: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, continuous outcome, n=800, $\sigma=200$, 5000 simulations.
+$ACE$ was approximately -60.
 Black squares denote the mean variance ratio (=SER).
-Note the $0.08\%$ of correct model specification simulations and $3.4\%$ of misspecified outcome model simulations where the ratio was above 1.2 are not displayed.
-Figure 4: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, n=800, $\sigma=200$, 5000 simulations.
+Results exclude 15 simulations where models failed to converge.
+The $0.14\%$ of correct model specification simulations, $4.5\%$ of misspecified outcome model simulations, and $0.06\%$ of misspecified propensity model simulations where the ratio was above 1.2 or below 0.8 are not displayed.
+Figure A3: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, continuous outcome, n=800, $\sigma=600$, 5000 simulations.
+$ACE$ was approximately -60.
 Black squares denote the mean variance ratio (=SER).
-Results exclude 14 simulations where models did not converge.
-Note the $0.06\%$ of correct model specification simulations, $5.2\%$ of misspecified outcome model simulations, and $0.01\%$ of simulations where both models were misspecified where the ratio was above 1.2 or below 0.8 are not displayed.
-Figure 5: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, n=2000, $\sigma=200$, 5000 simulations.
+The $0.77\%$ of correct model specification simulations, $3.66\%$ of misspecified outcome model simulations, and $0.01\%$ of misspecified propensity model simulations where the ratio was above 1.2 or below 0.8 are not displayed.
+Figure A4: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, continuous outcome, n=2000, $\sigma=200$, 5000 simulations.
+$ACE$ was approximately -60.
 Black squares denote the mean variance ratio (=SER).
-Note the $3.9\%$ of misspecified outcome model simulations where the ratio was above 1.2 are not displayed.
+The $2.9\%$ of misspecified outcome model simulations where the ratio was above 1.2 or below 0.8 are not displayed.
+Figure A5: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, continuous outcome, n=2000, $\sigma=400$, 5000 simulations.
+$ACE$ was approximately -60.
+Black squares denote the mean variance ratio (=SER).
+The $0.08\%$ of correct model specification simulations and $2.61\%$ of misspecified outcome model simulations where the ratio was above 1.2 or below 0.8 are not displayed.
+Figure A6: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, continuous outcome, n=2000, $\sigma=600$, 5000 simulations.
+$ACE$ was approximately -60.
+Black squares denote the mean variance ratio (=SER).
+The $0.19\%$ of correct model specification simulations and $2.33\%$ of misspecified outcome model simulations where the ratio was above 1.2 or below 0.8 are not displayed.
+Figure A7: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, continuous outcome, n=2000, $\sigma=400$, 5000 simulations under the null.
+Black squares denote the mean variance ratio (=SER).
+The $0.90\%$ of misspecified outcome model specification simulations where the ratio was above 2.75 or below 0.5 are not displayed.
+Figure A8: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, binary outcome, n=800, 5000 simulations.
+$ACE$ was approximately 0.045.
+Black squares denote the mean variance ratio (=SER).
+Results exclude eight simulations where models failed to converge.
+The $1.63\%$ of misspecified outcome model simulations and $0.01\%$ of simulations where both models were misspecified where the ratio was above 1.5 or below 0.8 are not displayed.
+Figure A9: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, binary outcome, n=800, 5000 simulations under the null.
+Black squares denote the mean variance ratio (=SER).
+Results exclude eight simulations where models failed to converge.
+The $0.27\%$ of correct model specification simulations, $2.26\%$ of misspecified outcome model simulations, $0.29\%$ of misspecified propensity simulations, and $0.01\%$ of simulations where both models were misspecified where the ratio was above 1.6 or below 0.8 are not displayed.
+Figure A10: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, binary outcome, n=2000, 5000 simulations.
+$ACE$ was approximately 0.045.
+Black squares denote the mean variance ratio (=SER).
+Results exclude four simulations where models failed to converge.
+The $0.72\%$ of misspecified outcome model simulations and $0.002\%$ of simulations where both models were misspecified where the ratio was above 1.5 or below 0.8 are not displayed.
+Figure A11: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, binary outcome, n=2000, 5000 simulations under the null.
+Black squares denote the mean variance ratio (=SER).
+Results exclude one simulation where models failed to converge.
+The $0.86\%$ of misspecified outcome model simulations and $0.002\%$ of simulations where both models were misspecified where the ratio was above 1.6 or below 0.8 are not displayed.